### v1

# Functions of bounded variation and Lipschitz algebras in metric measure spaces

Abstract.
Given a unital algebra $\mathscr{A}$ of locally Lipschitz functions defined over a metric measure space $({\rm X},{\sf d},\mathfrak{m})$, we study two associated notions of function of bounded variation and their relations: the space ${\rm BV}_{\rm H}({\rm X};\mathscr{A})$, obtained by approximating in energy with elements of $\mathscr{A}$, and the space ${\rm BV}_{\rm W}({\rm X};\mathscr{A})$, defined through an integration-by-parts formula that involves derivations acting in duality with $\mathscr{A}$. Our main result provides a sufficient condition on the algebra $\mathscr{A}$ under which ${\rm BV}_{\rm H}({\rm X};\mathscr{A})$ coincides with the standard metric BV space ${\rm BV}_{\rm H}({\rm X})$, which corresponds to taking as $\mathscr{A}$ the collection of all locally Lipschitz functions. Our result applies to several cases of interest, for example to Euclidean spaces and Riemannian manifolds equipped with the algebra of smooth functions, or to Banach and Wasserstein spaces equipped with the algebra of cylinder functions. Analogous results for metric Sobolev spaces ${\rm H}^{1,p}$ of exponent $p\in(1,\infty)$ were previously obtained by several different authors.
Key words and phrases:
Functions of bounded variation; Lipschitz algebras; metric measure spaces; derivations
2020 Mathematics Subject Classification:
53C23, 26A45, 49J52, 46E35, 46N10
1. Introduction
1.1. General overview
For more than two decades, functions of bounded variation and sets of finite perimeter have been studied in the general setting of metric measure spaces, starting from the paper [32]. Rather refined results are available on those metric measure spaces (often called PI spaces) that are doubling and support a weak Poincaré inequality (see e.g. [32, 1, 2, 25, 26, 27]), and even deeper structural properties have been proven when also a lower synthetic Ricci curvature bound is imposed, in the class of the so-called ${\sf RCD}(K,N)$ spaces [3, 13, 12, 11, 8]. However, a fully consistent theory of BV functions does not require any additional assumptions on the metric measure space under consideration, and that will be our object of study in the present paper.
The first notion of metric BV space, introduced by Miranda Jr. in [32], is formulated in terms of an energy approximation by locally Lipschitz functions. However, the algebra of locally Lipschitz functions depends solely on the metric structure of the ambient space, and as such it is not capable of detecting some additional features that the underlying metric space may have. Due to this reason, in several cases of interest – which we shall briefly mention in Section 1.3 – it is desirable to know that a smaller, specific algebra of approximating functions can be used to define the BV space. This is exactly the primary goal of the present paper: to give a sufficient (and effective) condition on some algebra of functions in order that the corresponding BV space (obtained via approximation) coincides with the original one that was defined in terms of locally Lipschitz functions.
More than ten years after the BV space via approximation was developed, Ambrosio and Di Marino introduced in [4] another notion of metric BV space and showed its equivalence with the former. Their definition is expressed in terms of the behaviour of functions along suitably-selected curves, where the exceptional curve families are detected using the so-called test plans. A third approach was then proposed by Di Marino in [17, 16], by means of an integration-by-parts formula involving an appropriate notion of derivation with divergence. Also the latter approach turned out to be equivalent to the first two. Generalised versions of the BV space via derivations will have a key role in this paper, as we shall discuss more in details later. Finally, we mention that a fourth notion of metric ‘Newtonian-type’ BV space (that we will not employ in this paper) was introduced more recently by Martio in [30, 31]. His notion is based upon the behaviour of functions along $AM$-almost every rectifiable curve, where $AM$ denotes the so-called approximation modulus. The full equivalence of this latter approach with the other ones we discussed above was proved in [33].
1.2. Contents of the paper
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space (as in Definition 2.1 below). For any locally Lipschitz function $f\colon{\rm X}\to\mathbb{R}$, we denote by ${\rm lip}_{a}(f)\colon{\rm X}\to[0,+\infty)$ its asymptotic slope function (see (2.1)), which assigns to each point $x\in{\rm X}$ the ‘infinitesimal Lipschitz constant’ of $f$ at $x$. The algebra of test functions we will consider in the various definitions of BV space is that of bounded locally Lipschitz functions whose asymptotic slope is integrable, which we denote by ${\rm LIP}_{\star}({\rm X})$; see Definition 2.2. Given any unital separating subalgebra $\mathscr{A}$ of ${\rm LIP}_{\star}({\rm X})$, we define:
  • The space ${\rm BV}_{\rm H}({\rm X};\mathscr{A})$ of those functions $f\in L^{1}(\mathfrak{m})$ that can be approximated in energy by elements of $\mathscr{A}$; see Definition 3.1. Any given $f\in{\rm BV}_{\rm H}({\rm X};\mathscr{A})$ is associated with a quantity $\|{\bf D}f\|_{*,\mathscr{A}}\in[0,+\infty)$, which we call its total variation (see (3.1)). In the distinguished case $\mathscr{A}={\rm LIP}_{\star}({\rm X})$, where we use the shorter notations ${\rm BV}_{\rm H}({\rm X})$ and $\|{\bf D}f\|_{*}$, our definition is consistent with the one of [32, 4]. Whereas each function $f\in{\rm BV}_{\rm H}({\rm X})$ is also naturally associated with a total variation measure $|{\bf D}f|_{*}$ satisfying $|{\bf D}f|_{*}({\rm X})=\|{\bf D}f\|_{*}$ (see Definition 3.2), we do not know whether it is possible to define some total variation measure associated to an arbitrary function $f\in{\rm BV}_{\rm H}({\rm X};\mathscr{A})$ when $\mathscr{A}\neq{\rm LIP}_{\star}({\rm X})$. Indeed, the fact that the set-valued function $|{\bf D}f|_{*}$ as in Definition 3.2 is actually a $\sigma$-additive measure is due to the fact that ${\rm LIP}_{\star}({\rm X})$ consists of a sufficiently vast class of locally Lipschitz functions, for which the results we will present in Appendix A are satisfied.
  • The space ${\rm BV}_{\rm W}({\rm X};\mathscr{A})$ of those functions $f\in L^{1}(\mathfrak{m})$ that satisfy an integration-by-parts formula in duality with the space ${\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ of Lipschitz $\mathscr{A}$-derivations $b\colon\mathscr{A}\to L^{1}(\mathfrak{m})$ having divergence ${\rm div}(b)\in L^{\infty}(\mathfrak{m})$; see Definitions 2.6, 2.7 and 3.10. Any $f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A})$ comes with a corresponding total $\mathscr{A}$-variation measure $|{\bf D}f|_{\mathscr{A}}$, as it is shown in Proposition 3.12. In the special case $\mathscr{A}={\rm LIP}_{\star}({\rm X})$, we just write ${\rm BV}_{\rm W}({\rm X})$ and $|{\bf D}f|$ for the sake of brevity. The original definition in [17, 16] corresponds to choosing as $\mathscr{A}$ the space of all boundedly-supported Lipschitz functions.
It is worth pointing out that – differently from most of the previous literature on the topic – in this paper we are not assuming completeness nor separability of the ambient metric space $({\rm X},{\sf d})$. The possible lack of separability does not cause major issues, since we still require the reference measure $\mathfrak{m}$ to be Radon (recall Definition 2.1). On the other hand, allowing for non-complete metric spaces leads to additional difficulties, as we will see in the sequel. We prefer to consider possibly non-complete spaces in order to include e.g. open domains in a given ambient space. Non-separable spaces are also interesting e.g. in view of the potential generalisation of the theories we consider here to the setting of extended metric-topological measure spaces [5, 36]. Let us also highlight that here we assume $\mathfrak{m}$ to be a finite measure. This choice was made mostly for convenience, as it allows us to work with unital subalgebras of ${\rm LIP}_{\star}({\rm X})$; this would not be possible for infinite reference measures, because constant functions are integrable only with respect to a finite measure. Extending our definitions and results to metric spaces equipped with a (possibly infinite) boundedly-finite Radon measure would be interesting, but outside the scope of this paper.
The main results of the present paper can be briefly summarised as follows:
  • For any metric measure space $({\rm X},{\sf d},\mathfrak{m})$ and any unital separating subalgebra $\mathscr{A}\subseteq{\rm LIP}_{\star}({\rm X})$, we prove in Theorem 4.8 that
    ${\rm BV}_{\rm W}({\rm X};\mathscr{A})\subseteq{\rm BV}_{\rm H}({\rm X};\mathscr{A}),$
    as well as the inequality $\|{\bf D}f\|_{*,\mathscr{A}}\leq|{\bf D}f|_{\mathscr{A}}({\rm X})$ for every $f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A})$. The verification of this statement is based on a rather general duality argument in Convex Analysis, which is inspired by the proof of [35, Theorem 5.4] (and [28, Theorem 3.3]), where metric Sobolev spaces are involved instead.
  • Assuming either that $({\rm X},{\sf d})$ is complete or that its associated topological space is a Radon space, we obtain in Theorem 4.9 the identification
    ${\rm BV}_{\rm H}({\rm X})={\rm BV}_{\rm W}({\rm X})$
    together with the identity of measures $|{\bf D}f|=|{\bf D}f|_{*}$ for every $f\in{\rm BV}_{\rm H}({\rm X})$. At this level of generality, this result seems to be new. In the case where $({\rm X},{\sf d})$ is complete, we can also deduce from the results of [4] that every function $f\in{\rm BV}_{\rm H}({\rm X})$ can be approximated in energy by a sequence of bounded globally Lipschitz functions; see Proposition 3.5. Here, the completeness assumption cannot be dropped, as we will showcase in Example 3.6.
  • Assuming that $({\rm X},{\sf d})$ is complete and $\mathscr{A}$ is a good algebra, we prove in Theorem 4.10 that
    ${\rm BV}_{\rm H}({\rm X};\mathscr{A})={\rm BV}_{\rm H}({\rm X}),$ (1.1)
    with $\|{\bf D}f\|_{*,\mathscr{A}}=\|{\bf D}f\|_{*}$ for every $f\in{\rm BV}_{\rm H}({\rm X})$. In other words, being a good algebra is a sufficient condition for $\mathscr{A}$ to be dense in energy in ${\rm BV}_{\rm H}({\rm X})$. By a ‘good algebra’ we mean a unital separating subalgebra $\mathscr{A}$ of ${\rm LIP}_{\star}({\rm X})$ consisting of bounded globally Lipschitz functions with the property that truncated distance functions from a point can be ‘approximated well in ${\rm H}^{1,1}$’ by elements of $\mathscr{A}$, where the Sobolev space ${\rm H}^{1,1}({\rm X};\mathscr{A})$ we consider is the one discussed in Definition 3.7; for details, we refer to Definition 4.1, where we will introduce the concept of good algebra. The proof of (1.1) relies on the key approximation result Proposition 4.4, where the assumption of $\mathscr{A}$ being a good algebra is used in an essential way. Along the way, we also obtain that ${\rm BV}_{\rm H}({\rm X};\mathscr{A})={\rm BV}_{\rm W}({\rm X};\mathscr{A})$.
1.3. Motivations and comparison with the previous literature
In studying functions of bounded variation defined using different algebras of locally Lipschitz functions, we have been strongly inspired by the works [36, 20], where unital separating subalgebras $\mathscr{A}$ of bounded Lipschitz functions were used to define metric Sobolev spaces ${\rm H}^{1,p}({\rm X};\mathscr{A})$ of exponent $p\in(1,\infty)$. It is shown in [36] that – in the more general framework of extended metric-topological measure spaces – if $\mathscr{A}$ is a compatible algebra (cf. with Example 4.2), then ${\rm H}^{1,p}({\rm X};\mathscr{A})$ coincides with the ‘standard’ Sobolev space ${\rm H}^{1,p}({\rm X})$ defined via approximation with bounded Lipschitz functions. Later on, a characterisation of those Lipschitz algebras $\mathscr{A}$ for which ${\rm H}^{1,p}({\rm X};\mathscr{A})={\rm H}^{1,p}({\rm X})$ has been provided in [20] (in the case where the ambient space is a metric measure space). Our main Theorem 4.10 can be regarded as a version of the above result for metric BV spaces, but with some differences:
  • We give only a sufficient condition for the identification ${\rm BV}_{\rm H}({\rm X};\mathscr{A})={\rm BV}_{\rm H}({\rm X})$, namely the fact that $\mathscr{A}$ is a good algebra. Conversely, one can easily realise (see Proposition 4.11) that a necessary condition for ${\rm BV}_{\rm H}({\rm X};\mathscr{A})={\rm BV}_{\rm H}({\rm X})$ to hold is that $\mathscr{A}$ is a weakly good subalgebra, which we define in a similar way as good algebras, but where the approximation of truncated distance functions is done in BV (as opposed to ${\rm H}^{1,1}$). A natural guess is that $\mathscr{A}$ being a weakly good subalgebra is in fact equivalent to ${\rm BV}_{\rm H}({\rm X};\mathscr{A})={\rm BV}_{\rm H}({\rm X})$. However, we were not able to prove it, and thus we leave it as an open problem. Another guess would be that $\mathscr{A}$ being a good algebra is equivalent to ${\rm H}^{1,1}({\rm X};\mathscr{A})={\rm H}^{1,1}({\rm X})$, but this problem is – arguably – even more difficult, due to the functional-analytic complications one encounters when studying metric Sobolev spaces of exponent $p=1$ (cf. with [7]).
  • Our proof strategy for Theorem 4.10 is different from the one of [36, 20]. In the latter works, the identification results are obtained directly at the level of the spaces ${\rm H}^{1,p}({\rm X};\mathscr{A})$, making use of (generalised) Hopf–Lax flow techniques. In this paper, instead, we first prove equivalence results for metric BV spaces defined using derivations, and then we show their equivalence with the BV spaces defined via approximation. As a theory of derivations on extended metric-topological measure spaces has been recently developed in [35], we expect that it is possible to extend many of the notions and results of this paper to that setting.
We conclude the introduction by discussing some motivations behind our interest in good algebras. As we will observe in Example 4.2, the family of good algebras includes all the compatible algebras considered by Savaré in [36], thus in particular the algebra of smooth functions in the Euclidean space (or in any Riemannian manifold), as well as the algebra of cylinder functions in a Banach space or in Wasserstein/Hellinger/Hellinger–Kantorovich spaces. As it is evident e.g. from the works [18, 36, 20, 37, 15], whenever the ambient space possesses ‘nice’ structural features, in order to investigate fine analytic properties it is of fundamental importance to detect a family of regular functions, which are associated with a ‘geometric’ notion of differential. We point out that Theorem 4.10 was already known to be true in two cases: for ${\rm X}=\mathbb{R}^{n}$ with the Euclidean norm, $\mathfrak{m}$ a Radon measure on $\mathbb{R}^{n}$ and $\mathscr{A}=C^{\infty}_{c}(\mathbb{R}^{n})$ the algebra of compactly-supported smooth functions, see [21, Theorem 5.7]; for ${\rm X}=\mathbb{B}$ a separable Banach space, $\mathfrak{m}$ a finite Borel measure on $\mathbb{B}$ and $\mathscr{A}={\rm Cyl}(\mathbb{B})$ the algebra of smooth cylindrical functions, see [34, Theorem 4.1].
Acknowledgements
The first named author was supported by the Research Council of Finland grant 362898.
2. Preliminaries
Let us begin by fixing some terminology and conventions. Throughout this paper, by an algebra we mean a commutative, associative algebra over the field $\mathbb{R}$ of real numbers. We say that an algebra $\mathscr{A}$ is unital if it has a multiplicative unit. Whenever we say that $\mathscr{A}^{\prime}$ is a unital subalgebra of $\mathscr{A}$, we implicitly make the assumption that $\mathscr{A}$ and $\mathscr{A}^{\prime}$ share the same multiplicative unit.
Given any real number $a\in\mathbb{R}$, we denote by $a^{+}\coloneqq a\vee 0$ and $a^{-}\coloneqq a\wedge 0$ its positive part and its negative part, respectively. Given a set ${\rm X}$, we indicate by $\mathbbm{1}_{S}$ the characteristic function of a subset $S$ of ${\rm X}$, i.e. we define $\mathbbm{1}_{S}(x)\coloneqq 1$ for every $x\in S$ and $\mathbbm{1}_{S}(x)\coloneqq 0$ for every $x\in{\rm X}\setminus S$.
For any two Banach spaces $\mathbb{B}_{1}$ and $\mathbb{B}_{2}$, we denote by $\mathcal{L}(\mathbb{B}_{1};\mathbb{B}_{2})$ the vector space of all bounded linear operators between $\mathbb{B}_{1}$ and $\mathbb{B}_{2}$. Recall that $\mathcal{L}(\mathbb{B}_{1};\mathbb{B}_{2})$ is a Banach space with respect to the operator norm, given by $\|T\|_{\mathcal{L}(\mathbb{B}_{1};\mathbb{B}_{2})}\coloneqq\sup\{\|T(v)\|_{\mathbb{B}_{2}}:v\in\mathbb{B}_{1},\,\|v\|_{\mathbb{B}_{1}}\leq 1\}$ for all $T\in\mathcal{L}(\mathbb{B}_{1};\mathbb{B}_{2})$. We denote by $\mathbb{B}^{\prime}:=\mathcal{L}(\mathbb{B};\mathbb{R})$ the topological dual of a Banach space $\mathbb{B}$.
2.1. Metric/measure spaces
In this section, we discuss several concepts and results concerning spaces that are equipped with a distance and/or a measure. Given a measure space $({\rm X},\Sigma,\mu)$ and an exponent $p\in[1,\infty]$, we denote by $(L^{p}(\mu),\|\cdot\|_{L^{p}(\mu)})$ the Lebesgue space of exponent $p$ over $({\rm X},\Sigma,\mu)$. We write $L^{p}(\mu)^{+}$ to indicate the set of all $f\in L^{p}(\mu)$ satisfying $f\geq 0$ $\mu$-a.e. on ${\rm X}$. We also recall that $L^{p}(\mu)$ is a Riesz space if endowed with the usual pointwise $\mu$-a.e. order. Under suitable assumptions (for example, if $\mu$ is $\sigma$-finite), it holds that $L^{p}(\mu)$ is a Dedekind complete lattice, which means that every order-bounded subset of $L^{p}(\mu)$ has a supremum and an infimum. For any set $\rm M$ of non-negative measures on $({\rm X},\Sigma)$, we denote its supremum measure by $\bigvee{\rm M}$:
$\bigvee{\rm M}(E)\coloneqq\sup\bigg{\{}\sum_{n\in\mathbb{N}}\mu_{n}(E_{n})\;\bigg{|}\;(\mu_{n})_{n\in\mathbb{N}}\subseteq{\rm M},\,(E_{n})_{n\in\mathbb{N}}\subseteq\Sigma\text{ partition of }E\bigg{\}}\quad\text{ for every }E\in\Sigma.$
We recall that $\bigvee{\rm M}$ is the least measure on $({\rm X},\Sigma)$ satisfying $\bigvee{\rm M}\geq\mu$ for every $\mu\in{\rm M}$, where the partial order $\leq$ on the set of all non-negative measures on $({\rm X},\Sigma)$ is defined in the following way: we declare that $\mu\leq\nu$ if and only if $\mu(E)\leq\nu(E)$ for every $E\in\Sigma$.
Let $({\rm X},{\sf d})$ be a given metric space. We denote by ${\rm C}({\rm X})$ the unital algebra (with unit $\mathbbm{1}_{\rm X}$) of all real-valued continuous functions on ${\rm X}$. We also consider its unital subalgebra ${\rm C}_{b}({\rm X})$ consisting of all bounded continuous functions. If $(\mu_{n})_{n\in\mathbb{N}}$ and $\mu$ are finite non-negative Borel measures on ${\rm X}$, then we say that $\mu_{n}$ weakly converges to $\mu$ if
$\int f\,{\mathrm{d}}\mu=\lim_{n\to\infty}\int f\,{\mathrm{d}}\mu_{n}\quad\text{ for every }f\in{\rm C}_{b}({\rm X}).$
We denote by $\mathcal{M}({\rm X})$ the vector space of all finite, signed Radon measures on ${\rm X}$. Then $\mathcal{M}({\rm X})$ is a Banach space if endowed with the total variation norm. We denote by $\mathcal{M}_{+}({\rm X})$ the set of all those measures $\mu\in\mathcal{M}({\rm X})$ that are non-negative. To any measure $\mu\in\mathcal{M}({\rm X})$, we associate its total variation measure $|\mu|\in\mathcal{M}_{+}({\rm X})$. The support of a measure $\mu\in\mathcal{M}({\rm X})$ is the closed subset ${\rm spt}(\mu)$ of ${\rm X}$ given by
${\rm spt}(\mu)\coloneqq\big{\{}x\in{\rm X}\;\big{|}\;|\mu|(B_{r}(x))>0\text{ for every }r>0\big{\}},$
where $B_{r}(x)\coloneqq\{y\in{\rm X}:{\sf d}(x,y)<r\}$ denotes the open ball in ${\rm X}$ of center $x$ and radius $r$. Since the inner regularity of $\mu$ ensures that $\mu$ is concentrated on a $\sigma$-compact set, it can be readily checked that ${\rm spt}(\mu)$ is separable and $\mu$ is concentrated on ${\rm spt}(\mu)$. With a slight abuse of terminology, we say that a metric space $({\rm X},{\sf d})$ is Radon if its induced topological space is a Radon space, i.e. every finite Borel measure defined on it is a Radon measure. Since all Souslin spaces are Radon spaces, we have that every (Borel subset of a) complete and separable metric space is Radon. See e.g. [9, Theorem 7.4.3, Section 7.14(vii)] and the references therein for a more detailed discussion.
A distinguished subalgebra of ${\rm C}({\rm X})$ is the space of all real-valued locally Lipschitz functions on ${\rm X}$, which is given by
${\rm LIP}_{loc}({\rm X})\coloneqq\bigg{\{}f\in{\rm C}({\rm X})\;\bigg{|}\;\inf_{r>0}{\rm Lip}(f;B_{r}(x))<+\infty\text{ for every }x\in{\rm X}\bigg{\}},$
where the Lipschitz constant ${\rm Lip}(f;E)$ of the function $f$ on a set $E\subseteq{\rm X}$ is defined as
${\rm Lip}(f;E)\coloneqq\sup\bigg{\{}\frac{|f(x)-f(y)|}{{\sf d}(x,y)}\;\bigg{|}\;x,y\in E,\,x\neq y\bigg{\}}\in[0,+\infty],$
where we adopt the convention that $\sup(\varnothing)\coloneqq 0$. The asymptotic slope ${\rm lip}_{a}(f)\colon{\rm X}\to[0,+\infty)$ of a function $f\in{\rm LIP}_{loc}({\rm X})$ is given by
${\rm lip}_{a}(f)(x)\coloneqq\inf_{r>0}{\rm Lip}(f;B_{r}(x))\quad\text{ for every }x\in{\rm X}.$ (2.1)
It is easy to check that ${\rm lip}_{a}(f)$ is upper semicontinuous (thus, Borel measurable). The space of all (globally) Lipschitz functions, i.e. of those $f\in{\rm LIP}_{loc}({\rm X})$ with ${\rm Lip}(f)\coloneqq{\rm Lip}(f;{\rm X})<+\infty$, is denoted by ${\rm LIP}({\rm X})$. Note that ${\rm LIP}_{b}({\rm X})\coloneqq{\rm C}_{b}({\rm X})\cap{\rm LIP}({\rm X})$ is a unital subalgebra of ${\rm C}_{b}({\rm X})$.
In this paper, we focus on the following class of metric measure spaces:
Definition 2.1 (Metric measure space).
We say that a triple $({\rm X},{\sf d},\mathfrak{m})$ is a metric measure space if $({\rm X},{\sf d})$ is a metric space and $\mathfrak{m}\in\mathcal{M}_{+}({\rm X})$ (i.e. $\mathfrak{m}\geq 0$ is a finite Radon measure on ${\rm X}$).
We point out that this notion of metric measure space is different from other approaches considered in the literature. In this paper, the following subalgebra of ${\rm C}_{b}({\rm X})$ will have a key role.
Definition 2.2 (The algebra ${\rm LIP}_{\star}({\rm X})$).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Then we define
${\rm LIP}_{\star}({\rm X})\coloneqq\big{\{}f\in{\rm C}_{b}({\rm X})\cap{\rm LIP}_{loc}({\rm X})\;\big{|}\;{\rm lip}_{a}(f)\in L^{1}(\mathfrak{m})\big{\}}.$
Note that ${\rm LIP}_{\star}({\rm X})$ depends both on the distance ${\sf d}$ and the measure $\mathfrak{m}$, even though we do not indicate it for brevity. Since ${\rm LIP}_{\star}({\rm X})$ is a vector subspace of ${\rm C}_{b}({\rm X})$ and $fg\in{\rm LIP}_{\star}({\rm X})$ for every $f,g\in{\rm LIP}_{\star}({\rm X})$ (as ${\rm lip}_{a}(fg)\leq|f|{\rm lip}_{a}(g)+|g|{\rm lip}_{a}(f)\in L^{1}(\mathfrak{m})$), we have that ${\rm LIP}_{\star}({\rm X})$ is a unital subalgebra of ${\rm C}_{b}({\rm X})$. We will work also with unital separating subalgebras $\mathscr{A}$ of ${\rm LIP}_{\star}({\rm X})$, i.e.
$\sup_{f\in\mathscr{A}}|f(x)-f(y)|>0\quad\text{ for every }x,y\in{\rm X}\text{ with }x\neq y.$
By the Stone–Weierstrass theorem, if $\mathscr{A}$ is a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$, then
[…]
see e.g. [36, Lemma 2.1.27] (notice that, even if there the algebra $\mathscr{A}$ is assumed to be compatible, one can easily check that the simple proof of the mentioned Lemma does not require additional assumptions on $\mathscr{A}$). It can be readily checked that ${\rm LIP}_{\star}({\rm X})$ itself is a separating algebra.
In the sequel, we will need a partition-of-unity-type result involving functions of a given unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$, see Lemma 2.4 below. To prove it, we will use the following simple consequence of Weierstrass’ approximation theorem:
Lemma 2.3.
Let $a,b\in\mathbb{R}$ with $a<b$ and let $f\in{\rm C}([a,b])$ be such that $f(t)>0$ for every $t\in[a,b]$. Then there exists a sequence $(p_{n})_{n}$ of polynomials such that $0\leq p_{n}(t)\leq f(t)$ for every $t\in[a,b]$ and every $n\in\mathbb{N}$, and $p_{n}$ converges uniformly on $[a,b]$ to $f$ as $n\to\infty$.
Proof.
Let $\alpha:=\min_{[a,b]}f>0$. For every $n\in\mathbb{N}$ with $n>n_{0}\coloneqq\lceil 2/\alpha\rceil$, by the Weierstrass approximation theorem, there is a polynomial $\tilde{p}_{n}$ such that $\|f-\tilde{p}_{n}\|_{{\rm C}_{b}([a,b])}<1/n$. It is then enough to set $p_{n}\coloneqq\tilde{p}_{n_{0}+n}-\frac{1}{n+n_{0}}$ for every $n\in\mathbb{N}$. ∎
Lemma 2.4.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space, $\mathscr{A}$ a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$ and $E_{1},\ldots,E_{n}$ a Borel partition of ${\rm X}$. Fix any $\varepsilon>0$. Then there exist $\eta_{1},\ldots,\eta_{n}\in\mathscr{A}$ such that the following properties hold:
  • $\rm i)$
    $0\leq\eta_{i}\leq 1$ for every $i=1,\ldots,n$.
  • $\rm ii)$
    $\sum_{i=1}^{n}\eta_{i}(x)\leq 1$ for every $x\in{\rm X}$.
  • $\rm iii)$
    $\|\eta_{i}-\mathbbm{1}_{E_{i}}\|_{L^{1}(\mathfrak{m})}\leq\varepsilon$ for every $i=1,\ldots,n$.
Proof.
Fix $\delta>0$. Since $\mathscr{A}$ is dense in $L^{1}(\mathfrak{m})$, for any $i=1,\ldots,k$ we can find $(\psi^{k}_{i})_{k}\subseteq\mathscr{A}$ such that $\mathbbm{1}_{E_{i}}(x)=\lim_{k}\psi^{k}_{i}(x)$ for $\mathfrak{m}$-a.e. $x\in{\rm X}$. Thanks to [36, Corollary 2.1.24], we can also assume that $0\leq\psi^{k}_{i}\leq 1$. By Lemma 2.3, there is a polynomial $p$ such that
$0\leq p(t)\leq\frac{1}{t},\qquad\bigg{|}p(t)-\frac{1}{t}\bigg{|}\leq\delta\quad\text{ for every }t\in[\delta,n+\delta].$
Since $p\big{(}\sum_{j=1}^{n}\psi^{k}_{j}(x)+\delta\big{)}\to p(1+\delta)$ as $k\to\infty$ for $\mathfrak{m}$-a.e. $x\in{\rm X}$ and $p\circ\big{(}\sum_{j=1}^{n}\psi^{k}_{j}+\delta\big{)}\leq\frac{1}{\delta}$ holds $\mathfrak{m}$-a.e. on ${\rm X}$ for every $k\in\mathbb{N}$, by the dominated convergence theorem we can find $k_{0}\in\mathbb{N}$ such that, letting $\psi_{i}\coloneqq\psi^{k_{0}}_{i}$, it holds that $\big{\|}p\big{(}\sum_{j=1}^{n}\psi_{j}+\delta\big{)}-p(1+\delta)\big{\|}_{L^{1}(\mathfrak{m})}\leq\delta$ and $\|\psi_{i}-\mathbbm{1}_{E_{i}}\|_{L^{1}(\mathfrak{m})}\leq\delta$ for every $i=1,\ldots,n$. Now, let us define
$\eta_{i}\coloneqq p\bigg{(}\sum_{j=1}^{n}\psi_{j}+\delta\bigg{)}\psi_{i}\in\mathscr{A}\quad\text{ for every }i=1,\ldots,n.$
Notice that $\eta_{i}\geq 0$ for every $i=1,\ldots,n$ and $\sum_{i=1}^{n}\eta_{i}\leq\big{(}\sum_{j=1}^{n}\psi_{j}+\delta\big{)}^{-1}\big{(}\sum_{i=1}^{n}\psi_{i}\big{)}\leq 1$, which proves i) and ii). Finally, for any given $i=1,\ldots,n$ we can estimate
$\begin{split}\|\eta_{i}-\mathbbm{1}_{E_{i}}\|_{L^{1}(\mathfrak{m})}&\leq\bigg{\|}p\bigg{(}\sum_{j=1}^{n}\psi_{j}+\delta\bigg{)}\psi_{i}-\psi_{i}\bigg{\|}_{L^{1}(\mathfrak{m})}+\|\psi_{i}-\mathbbm{1}_{E_{i}}\|_{L^{1}(\mathfrak{m})}\\ &\leq\bigg{\|}p\bigg{(}\sum_{j=1}^{n}\psi_{j}+\delta\bigg{)}-\frac{1}{1+\delta}\bigg{\|}_{L^{1}(\mathfrak{m})}+\bigg{(}1-\frac{1}{1+\delta}\bigg{)}\mathfrak{m}({\rm X})+\delta\\ &\leq\bigg{\|}p\bigg{(}\sum_{j=1}^{n}\psi_{j}+\delta\bigg{)}-p(1+\delta)\bigg{\|}_{L^{1}(\mathfrak{m})}+\bigg{|}p(1+\delta)-\frac{1}{1+\delta}\bigg{|}\mathfrak{m}({\rm X})+(\mathfrak{m}({\rm X})+1)\delta\\ &\leq 2(\mathfrak{m}({\rm X})+1)\delta,\end{split}$
which proves the validity of iii) if we choose $\delta>0$ so that $2(\mathfrak{m}({\rm X})+1)\delta\leq\varepsilon$. ∎
2.2. $L^{p}$-Banach $L^{\infty}$-modules
Let us recall the language of $L^{p}$-Banach $L^{\infty}$-modules that was introduced in [23] (see also [22]). Following [23, Definition 1.2.10] (and with a slight change of terminology), we give the following definition:
Definition 2.5 ($L^{p}$-Banach $L^{\infty}$-module).
Let $({\rm X},\Sigma,\mu)$ be a $\sigma$-finite measure space and $p\in[1,\infty]$. Let $\mathscr{M}$ be a module over the commutative ring $L^{\infty}(\mu)$. Then we say that $\mathscr{M}$ is an $L^{p}(\mu)$-Banach $L^{\infty}(\mu)$-module if it is endowed with a map $|\cdot|\colon\mathscr{M}\to L^{p}(\mu)^{+}$, called pointwise norm, such that the following conditions are satisfied:
  • $\rm i)$
    Given any $v,w\in\mathscr{M}$ and $f\in L^{\infty}(\mu)$, it holds that
    $\begin{split}&|v|=0\quad\Longrightarrow\quad v=0,\\ &|v+w|\leq|v|+|w|,\\ &|fv|=|f||v|,\end{split}$
    where all equalities and inequalities involving the pointwise norm are in the $\mu$-a.e. sense.
  • $\rm ii)$
    The glueing property holds, i.e. if $(E_{n})_{n\in\mathbb{N}}\subseteq\Sigma$ is a partition of the set ${\rm X}$ and a sequence $(v_{n})_{n\in\mathbb{N}}\subseteq\mathscr{M}$ satisfies $\sum_{n\in\mathbb{N}}\|\mathbbm{1}_{E_{n}}|v_{n}|\|_{L^{p}(\mu)}^{p}<+\infty$, then there exists a (necessarily unique) element $v=\sum_{n\in\mathbb{N}}\mathbbm{1}_{E_{n}}v_{n}\in\mathscr{M}$ such that $\mathbbm{1}_{E_{n}}v=\mathbbm{1}_{E_{n}}v_{n}$ for every $n\in\mathbb{N}$.
  • $\rm iii)$
    The norm $\|v\|_{\mathscr{M}}\coloneqq\||v|\|_{L^{p}(\mu)}$ on $\mathscr{M}$ is complete.
Observe that each $L^{p}(\mu)$-Banach $L^{\infty}(\mu)$-module is in particular a Banach space. We say that a subset $S$ of $\mathscr{M}$ generates $\mathscr{M}$ provided the $L^{\infty}(\mu)$-linear span of $S$ is dense in $\mathscr{M}$. An isomorphism $\Phi\colon\mathscr{M}\to\mathscr{N}$ between two $L^{p}(\mu)$-Banach $L^{\infty}(\mu)$-modules $\mathscr{M}$ and $\mathscr{N}$ is an $L^{\infty}(\mu)$-linear bijection satisfying $|\Phi(v)|=|v|$ for every $v\in\mathscr{M}$.
Since any given $L^{1}(\mu)$-Banach $L^{\infty}(\mu)$-module $\mathscr{M}$ is a Banach space, we can consider its dual Banach space $\mathscr{M}^{\prime}$. Alternatively, we can consider its dual in the sense of Banach modules: following [23, Definition 1.2.6], we define $\mathscr{M}^{*}$ as the set of all $L^{\infty}(\mu)$-linear maps $\omega\colon\mathscr{M}\to L^{1}(\mu)$ for which there exists a function $g\in L^{\infty}(\mu)^{+}$ such that
$|\omega(v)|\leq g|v|\text{ holds }\mu\text{-a.e.\ on }{\rm X}\text{ for every }v\in\mathscr{M}.$ (2.3)
Then $\mathscr{M}^{*}$ is an $L^{\infty}(\mu)$-Banach $L^{\infty}(\mu)$-module if endowed with the pointwise operations and with
[…]
We say that $\mathscr{M}^{*}$ is the module dual of $\mathscr{M}$. The relation between $\mathscr{M}^{\prime}$ and $\mathscr{M}^{*}$ is clarified by the following result: letting $\textsc{Int}_{\mathscr{M}}\colon\mathscr{M}^{*}\to\mathscr{M}^{\prime}$ be the operator given by
$\textsc{Int}_{\mathscr{M}}(\omega)(v)\coloneqq\int\omega(v)\,{\mathrm{d}}\mathfrak{m}\quad\text{ for every }\omega\in\mathscr{M}^{*}\text{ and }v\in\mathscr{M},$
we know from [23, Proposition 1.2.13] that
[…]
2.3. Lipschitz derivations
In this paper, we consider Lipschitz derivations in the sense of [17, 16]. Borrowing from [17, Section 1.1] and [7, Section 4.1], we give the ensuing definitions:
Definition 2.6 (Derivation).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. Then a map $b\colon\mathscr{A}\to L^{1}(\mathfrak{m})$ is called a Lipschitz $\mathscr{A}$-derivation if:
  • $\rm i)$
    There exists $G\in L^{\infty}(\mathfrak{m})^{+}$ such that $|b(f)|\leq G\,{\rm lip}_{a}(f)$ holds $\mathfrak{m}$-a.e. for every $f\in\mathscr{A}$.
  • $\rm ii)$
    The map $b$ is a linear operator that satisfies the Leibniz rule, i.e.
    $b(fg)=f\,b(g)+g\,b(f)\quad\text{ for every }f,g\in\mathscr{A}.$
We denote by ${\rm Der}^{\infty}({\rm X};\mathscr{A})$ the space of all Lipschitz $\mathscr{A}$-derivations on $({\rm X},{\sf d},\mathfrak{m})$.
The space ${\rm Der}^{\infty}({\rm X};\mathscr{A})$ is a module over the ring $L^{\infty}(\mathfrak{m})$, thus in particular it is a vector space. Given any $b\in{\rm Der}^{\infty}({\rm X};\mathscr{A})$, we denote by $|b|_{\mathscr{A}}\in L^{\infty}(\mathfrak{m})^{+}$ the $\mathfrak{m}$-a.e. minimal function $G$ as in Definition 2.6 i), whose existence is guaranteed by the Dedekind completeness of $L^{\infty}(\mathfrak{m})$. The space $({\rm Der}^{\infty}({\rm X};\mathscr{A}),|\cdot|_{\mathscr{A}})$ is an $L^{\infty}(\mathfrak{m})$-Banach $L^{\infty}(\mathfrak{m})$-module. When $\mathscr{A}={\rm LIP}_{\star}({\rm X})$, we just write ${\rm Der}^{\infty}({\rm X})$ and $|b|$ for brevity.
Definition 2.7 (Divergence).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. Then we say that $b\in{\rm Der}^{\infty}({\rm X};\mathscr{A})$ has divergence ${\rm div}(b)\in L^{\infty}(\mathfrak{m})$ if
$\int b(f)\,{\mathrm{d}}\mathfrak{m}=-\int f\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m}\quad\text{ for every }f\in\mathscr{A}.$
We denote by ${\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ the space of all derivations $b\in{\rm Der}^{\infty}({\rm X};\mathscr{A})$ having divergence.
Remark 2.8 (Independence of ${\rm div}(b)$ from $\mathscr{A}$).
As $\mathscr{A}$ is dense in $L^{1}(\mathfrak{m})$ by (2.2), we have that ${\rm div}(b)$ is uniquely determined. In principle, ${\rm div}(b)$ may depend on the choice of the algebra $\mathscr{A}$. However, it is easy to check that if $\mathscr{A}\subseteq\tilde{\mathscr{A}}$ are unital separating subalgebras of ${\rm LIP}_{\star}({\rm X})$ and $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\tilde{\mathscr{A}})$, then $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ and the two notions of divergence (in duality with $\mathscr{A}$ and $\tilde{\mathscr{A}}$, respectively) coincide. Thus, with a slight abuse of notation, we will write only ${\rm div}(b)$ without stressing the dependence on the chosen algebra, since it will be clear from the context. $\blacksquare$
Note also that the space ${\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ is a module over the ring $\mathscr{A}$, thus in particular it is a vector subspace of ${\rm Der}^{\infty}({\rm X};\mathscr{A})$. We denote by
$\overline{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$
the closure of ${\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ in ${\rm Der}^{\infty}({\rm X};\mathscr{A})$. Note that $\overline{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ is a Banach space with respect to the norm induced by ${\rm Der}^{\infty}({\rm X};\mathscr{A})$.
3. Metric $\rm BV$ and ${\rm H}^{1,1}$ spaces
3.1. The space ${\rm BV}_{\rm H}$
A metric BV space defined in terms of approximating (locally) Lipschitz functions was introduced in [32] and revisited in [4]. Here, we consider its following variant:
Definition 3.1 (BV space via relaxation).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. Then we define ${\rm BV}_{\rm H}({\rm X};\mathscr{A})$ as the space of all $f\in L^{1}(\mathfrak{m})$ for which there exists $(f_{n})_{n}\subseteq\mathscr{A}$ such that $f_{n}\to f$ in $L^{1}(\mathfrak{m})$ and $\sup_{n\in\mathbb{N}}\int{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}<+\infty$.
The total variation of a function $f\in{\rm BV}_{\rm H}({\rm X};\mathscr{A})$ is given by the quantity
$\|{\bf D}f\|_{*,\mathscr{A}}\coloneqq\inf\bigg{\{}\liminf_{n\to\infty}\int{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}\;\bigg{|}\;(f_{n})_{n}\subseteq\mathscr{A},\,f_{n}\to f\text{ in }L^{1}(\mathfrak{m})\bigg{\}}<+\infty.$ (3.1)
Clearly, there is a sequence $(f_{n})_{n}\subseteq\mathscr{A}$ such that $f_{n}\to f$ in $L^{1}(\mathfrak{m})$ and $\int{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}\to\|{\bf D}f\|_{*,\mathscr{A}}$. The Cheeger $\mathscr{A}$-energy functional $\mathcal{E}_{\mathscr{A}}\colon L^{1}(\mathfrak{m})\to[0,+\infty]$ is then defined as
$\mathcal{E}_{\mathscr{A}}(f)\coloneqq\left\{\begin{array}[]{ll}\|{\bf D}f\|_{*,\mathscr{A}}\\ +\infty\end{array}\quad\begin{array}[]{ll}\text{ if }f\in{\rm BV}_{\rm H}({\rm X};\mathscr{A}),\\ \text{ otherwise.}\end{array}\right.$
It is a direct consequence of the definition that
${\rm BV}_{\rm H}({\rm X};\mathscr{A})\subseteq{\rm BV}_{\rm H}({\rm X};\tilde{\mathscr{A}}),\qquad\mathcal{E}_{\tilde{\mathscr{A}}}\leq\mathcal{E}_{\mathscr{A}}\quad\text{ whenever }\mathscr{A}\subseteq\tilde{\mathscr{A}}.$ (3.2)
In the case $\mathscr{A}={\rm LIP}_{\star}({\rm X})$, we use the shorthand notations ${\rm BV}_{\rm H}({\rm X})$ and $\|{\bf D}f\|_{*}$. To any given function in ${\rm BV}_{\rm H}({\rm X})$, one can associate also a total variation measure, as follows:
Definition 3.2 (Total variation measure $|{\bf D}f|_{*}$).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $f\in{\rm BV}_{\rm H}({\rm X})$ be given. Then for any open set $U\subseteq{\rm X}$ we define
$|{\bf D}f|_{*}(U)\coloneqq\inf\bigg{\{}\liminf_{n\to\infty}\int_{U}{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}\;\bigg{|}\;{\rm LIP}_{\star}(U)\ni f_{n}\to f|_{U}\text{ in }L^{1}(\mathfrak{m}|_{U})\bigg{\}}\in[0,\|{\bf D}f\|_{*}].$
Observe that $|{\bf D}f|_{*}({\rm X})=\|{\bf D}f\|_{*}$. Furthermore, by arguing as in [4, Lemma 5.2] one can show that the set-function introduced in Definition 3.2 induces via Carathéodory’s construction a finite Borel measure, still denoted by $|{\bf D}f|_{*}$. Videlicet, the Borel measure $|{\bf D}f|_{*}$ is characterised by
$|{\bf D}f|_{*}(E)\coloneqq\inf\big{\{}|{\bf D}f|_{*}(U)\;\big{|}\;U\subseteq{\rm X}\text{ open, }E\subseteq U\big{\}}\quad\text{ for every Borel set }E\subseteq{\rm X}.$
This is shown in Appendix A, reporting a detailed version of the arguments in the proofs of [4, Lemmata 5.2 and 5.4] showing that they do not make use of the completeness nor separability of the ambient metric space. Moreover, even though in [4, Eq. (5.1)] the approximating sequence in the definition of $|{\bf D}f|_{*}(U)$ is not required to consist of bounded functions, it still coincides with our definition thanks to a standard truncation argument.
Remark 3.3.
Assume $f\in{\rm BV}_{\rm H}({\rm X})$ and $(f_{n})_{n}\subseteq{\rm LIP}_{\star}({\rm X})$ are chosen so that $f_{n}\to f$ in $L^{1}(\mathfrak{m})$ and $\int{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}\to\|{\bf D}f\|_{*}$. Then it holds that
${\rm lip}_{a}(f_{n})\mathfrak{m}\rightharpoonup|{\bf D}f|_{*}\quad\text{ weakly.}$.
Indeed, for any $U\subseteq{\rm X}$ open we have $|{\bf D}f|_{*}(U)\leq\liminf_{n}\int_{U}{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}=\liminf_{n}({\rm lip}_{a}(f_{n})\mathfrak{m})(U)$. Taking into account also that $({\rm lip}_{a}(f_{n})\mathfrak{m})({\rm X})\to|{\bf D}f|_{*}({\rm X})$, we conclude that ${\rm lip}_{a}(f_{n})\mathfrak{m}\rightharpoonup|{\bf D}f|_{*}$ weakly, thanks to the Portmanteau theorem. $\blacksquare$
Remark 3.4.
It follows from Remark 3.3 that
$|{\bf D}f|_{*}\text{ is concentrated on }{\rm spt}(\mathfrak{m})\text{ for every }f\in{\rm BV}_{\rm H}({\rm X}).$
Indeed, as ${\rm X}\setminus{\rm spt}(\mathfrak{m})$ is open, we have that $|{\bf D}f|_{*}({\rm X}\setminus{\rm spt}(\mathfrak{m}))\leq\liminf_{n}\int_{{\rm X}\setminus{\rm spt}(\mathfrak{m})}{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}=0$. However, even though ${\rm spt}(\mathfrak{m})$ is separable, it could happen that $|{\bf D}f|_{*}$ is not a Radon measure, but under the additional assumption that $({\rm X},{\sf d})$ is complete it holds that $|{\bf D}f|_{*}$ is Radon (since in this case ${\rm spt}(\mathfrak{m})$ is complete and separable, thus it is a Radon space). $\blacksquare$
It is highly non-trivial to prove that, assuming the ambient metric space is complete, every function in ${\rm BV}_{\rm H}({\rm X})$ can be approximated (on the whole ${\rm X}$) by a sequence of bounded Lipschitz functions. The validity of this claim follows from the results of [16], as we are going to see:
Proposition 3.5.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space such that $({\rm X},{\sf d})$ is complete. Then
${\rm BV}_{\rm H}({\rm X})={\rm BV}_{\rm H}({\rm X};{\rm LIP}_{b}({\rm X}))$
and it holds that $\|{\bf D}f\|_{*,{\rm LIP}_{b}({\rm X})}=\|{\bf D}f\|_{*}$ for every $f\in{\rm BV}_{\rm H}({\rm X})$.
Proof.
Under the additional assumption that $({\rm X},{\sf d})$ is separable, the statement is a direct consequence of [16, Theorem 4.5.3]. Let us now show how to drop the separability assumption. For brevity, let us denote by $S$ the support of $\mathfrak{m}$. Defining the so-called weak-BV spaces ${\rm BV}_{w}$ as it is done in [16, Section 4.4.3] (or [4, Section 5.3]), by arguing as in [16, Section 4.4] one can check that
${\rm BV}_{\rm H}({\rm X};{\rm LIP}_{b}({\rm X}))\subseteq{\rm BV}_{\rm H}({\rm X})\subseteq{\rm BV}_{w}({\rm X}),$
as well as $|{\bf D}f|_{w}({\rm X})\leq\|{\bf D}f\|_{*}\leq\|{\bf D}f\|_{*,{\rm LIP}_{b}({\rm X})}$ for every $f\in{\rm BV}_{\rm H}({\rm X};{\rm LIP}_{b}({\rm X}))$, where $|{\bf D}f|_{w}$ denotes the total variation measure corresponding to the weak-BV space ${\rm BV}_{w}({\rm X})$. Observe that ${\rm BV}_{\rm H}({\rm X};{\rm LIP}_{b}({\rm X}))$ can be identified with ${\rm BV}_{\rm H}(S;{\rm LIP}_{b}(S))$ (thanks to [19]), and that ${\rm BV}_{w}({\rm X})$ can be identified with ${\rm BV}_{w}(S)$ (since $\infty$-test plans are concentrated on curves that lie in the support of $\mathfrak{m}$); in both cases, also the total variations coincide. Since ${\rm BV}_{\rm H}(S;{\rm LIP}_{b}(S))={\rm BV}_{w}(S)$ and $|{\bf D}f|_{w}(S)=\|{\bf D}f\|_{*,{\rm LIP}_{b}(S)}$ for every $f\in{\rm BV}_{w}(S)$ by [16, Theorem 4.5.3] (together with the fact that $S$ is complete and separable), we can finally conclude that the statement holds. ∎
We point out that the completeness assumption in Proposition 3.5 cannot be dropped, as it is shown by the following example:
Example 3.6.
Let us consider the space ${\rm X}\subseteq\mathbb{R}$, which we define as
${\rm X}\coloneqq\bigcup_{n=1}^{\infty}(n-2^{-n},n+2^{-n})\setminus\{n\}.$
We equip ${\rm X}$ with the restriction ${\sf d}$ of the Euclidean distance and the restriction $\mathfrak{m}$ of the one-dimensional Lebesgue measure. Note that $({\rm X},{\sf d},\mathfrak{m})$ is a metric measure space with $({\rm X},{\sf d})$ separable but non-complete. Let us now consider the function
$f\coloneqq\mathbbm{1}_{\bigcup_{n=1}^{\infty}(n,n+2^{-n})}\in{\rm LIP}_{\star}({\rm X}).$
Since ${\rm lip}_{a}(f)\equiv 0$ we see that $f\in{\rm BV}_{\rm H}({\rm X})$ with $\|{\bf D}f\|_{*}=0$. However, we claim that $f\notin{\rm BV}_{\rm H}({\rm X};{\rm LIP}_{b}({\rm X}))$. To prove it, fix an arbitrary sequence $(f_{j})_{j}\subseteq{\rm LIP}_{b}({\rm X})$ such that $f_{j}\to f$ in $L^{1}(\mathfrak{m})$. Note that each $f_{j}$ can be uniquely extended to a bounded Lipschitz function on $\bigcup_{n=1}^{\infty}[n-2^{-n},n+2^{-n}]$, which we still denote by $f_{j}$. Up to a passing to a non-relabelled subsequence, we can assume that $\liminf_{j}\int{\rm lip}_{a}(f_{j})\,{\mathrm{d}}\mathfrak{m}$ is in fact a limit. Up to a further subsequence, we also have that $f_{j}\to f$ in the pointwise $\mathfrak{m}$-a.e. sense. In particular, for any given $k\in\mathbb{N}$ we can find $j_{k}\in\mathbb{N}$ and $(t^{k}_{\ell})_{\ell=1}^{k},(s^{k}_{\ell})_{\ell=1}^{k}\subseteq{\rm X}$, with $\ell-2^{-\ell}<t^{k}_{\ell}<\ell<s^{k}_{\ell}<\ell+2^{-\ell}$ for all $\ell=1,\ldots,k$, such that $f_{j}(t^{k}_{\ell})\leq\frac{1}{3}$ and $f_{j}(s^{k}_{\ell})\geq\frac{2}{3}$ for every $\ell=1,\ldots,k$ and $j\geq j_{k}$. Then
$\frac{1}{3}\leq f_{j}(s^{k}_{\ell})-f_{j}(t^{k}_{\ell})=\int_{t^{k}_{\ell}}^{s^{k}_{\ell}}f^{\prime}_{j}(r)\,{\mathrm{d}}r\leq\int_{\ell-2^{-\ell}}^{\ell+2^{-\ell}}{\rm lip}_{a}(f_{j})(r)\,{\mathrm{d}}r\quad\text{ for all }\ell=1,\ldots,k\text{ and }j\geq j_{k}.$
Summing over $\ell=1,\ldots,k$ and letting $j\to\infty$, we deduce that
$\lim_{j\to\infty}\int{\rm lip}_{a}(f_{j})\,{\mathrm{d}}\mathfrak{m}\geq\limsup_{j\to\infty}\sum_{\ell=1}^{k}\int_{\ell-2^{-\ell}}^{\ell+2^{-\ell}}{\rm lip}_{a}(f_{j})(r)\,{\mathrm{d}}r\geq\frac{k}{3}.$
Due to the arbitrariness of $k\in\mathbb{N}$, we conclude that $\lim_{j}\int{\rm lip}_{a}(f_{j})\,{\mathrm{d}}\mathfrak{m}=+\infty$ and thus accordingly $f\notin{\rm BV}_{\rm H}({\rm X};{\rm LIP}_{b}({\rm X}))$. $\blacksquare$
3.2. The space ${\rm H}^{1,1}$
A metric Sobolev space defined in terms of approximating Lipschitz functions was introduced in [14] and revisited in [6]. Here, we consider its following variant:
Definition 3.7 (The space ${\rm H}^{1,1}$).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. Then we define ${\rm H}^{1,1}({\rm X};\mathscr{A})$ as the space of all $f\in L^{1}(\mathfrak{m})$ for which there exist a sequence $(f_{n})_{n}\subseteq\mathscr{A}$ and a function $G\in L^{1}(\mathfrak{m})^{+}$ such that $f_{n}\to f$ strongly in $L^{1}(\mathfrak{m})$ and ${\rm lip}_{a}(f_{n})\rightharpoonup G$ weakly in $L^{1}(\mathfrak{m})$.
Given any $f\in{\rm H}^{1,1}({\rm X};\mathscr{A})$, we say that $\tilde{G}\in L^{1}(\mathfrak{m})^{+}$ is a relaxed $\mathscr{A}$-slope of $f$ if $\tilde{G}\geq G$ for some $G\in L^{1}(\mathfrak{m})^{+}$ as in Definition 3.7. The set of all relaxed $\mathscr{A}$-slopes of $f$ is a closed convex sublattice of $L^{1}(\mathfrak{m})$, thus it admits a unique minimal element $|{\rm D}f|_{*,\mathscr{A}}$, which we call the minimal relaxed $\mathscr{A}$-slope of $f$. The space ${\rm H}^{1,1}({\rm X};\mathscr{A})$ is a Banach space if endowed with the norm
$\|f\|_{{\rm H}^{1,1}({\rm X};\mathscr{A})}\coloneqq\|f\|_{L^{1}(\mathfrak{m})}+\||{\rm D}f|_{*,\mathscr{A}}\|_{L^{1}(\mathfrak{m})}\quad\text{ for every }f\in{\rm H}^{1,1}({\rm X};\mathscr{A}).$
Observe that $\mathscr{A}\subseteq{\rm H}^{1,1}({\rm X};\mathscr{A})$ and
$|{\rm D}f|_{*,\mathscr{A}}\leq{\rm lip}_{a}(f)\quad\text{ for every }f\in\mathscr{A}.$ (3.3)
Minimal relaxed $\mathscr{A}$-slopes satisfy also the following locality property: for any $f,g\in{\rm H}^{1,1}({\rm X};\mathscr{A})$,
$|{\rm D}f|_{*,\mathscr{A}}=|{\rm D}g|_{*,\mathscr{A}}\quad\text{ holds }\mathfrak{m}\text{-a.e.\ on }\{f=g\}.$ (3.4)
The interested reader can find a proof of the existence of the minimal relaxed $\mathscr{A}$-slope and of the completeness of $H^{1,p}({\rm X};\mathscr{A})$ in [7, Section 5.2] for $\mathscr{A}$ being the algebra of Lipschitz functions with bounded support, or in [36, Section 3.1] in case $p>1$. The locality property of the relaxed $\mathscr{A}$-slope can be proven exactly as in [36, Theorem 3.1.12(b)], where it is proven for $p>1$.
In the case where $\mathscr{A}={\rm LIP}_{\star}({\rm X})$, we use the shorthand notations ${\rm H}^{1,1}({\rm X})$ and $|{\rm D}f|_{*}$.
Remark 3.8.
It holds that ${\rm H}^{1,1}({\rm X};\mathscr{A})\subseteq{\rm BV}_{\rm H}({\rm X};\mathscr{A})$ and
$\|{\bf D}f\|_{*,\mathscr{A}}\leq\int|{\rm D}f|_{*,\mathscr{A}}\,{\mathrm{d}}\mathfrak{m}\quad\text{ for every }f\in{\rm H}^{1,1}({\rm X};\mathscr{A}).$
The proof follows immediately from the definitions of the spaces and the weak lower semicontinuity of the $L^{1}(\mathfrak{m})$-norm. $\blacksquare$
In accordance with [23, Definition 2.2.1], the metric Sobolev space ${\rm H}^{1,1}({\rm X};\mathscr{A})$ induces a space of ‘integrable $1$-forms’, which is an $L^{1}(\mathfrak{m})$-Banach $L^{\infty}(\mathfrak{m})$-module called the cotangent module, and a differential map that underlies the minimal relaxed slopes. Videlicet:
Theorem 3.9 (Cotangent module).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. Then there exist an $L^{1}(\mathfrak{m})$-Banach $L^{\infty}(\mathfrak{m})$-module $L^{1}(T^{*}{\rm X};\mathscr{A})$, which we call the $\mathscr{A}$-cotangent module, and a linear operator ${\mathrm{d}}_{\mathscr{A}}\colon\mathscr{A}\to L^{1}(T^{*}{\rm X};\mathscr{A})$ such that:
  • $\rm i)$
    $|{\mathrm{d}}_{\mathscr{A}}f|=|{\rm D}f|_{*,\mathscr{A}}$ for every $f\in\mathscr{A}$.
  • $\rm ii)$
    $\{{\mathrm{d}}_{\mathscr{A}}f:f\in\mathscr{A}\}$ generates $L^{1}(T^{*}{\rm X};\mathscr{A})$.
The pair $(L^{1}(T^{*}{\rm X};\mathscr{A}),{\mathrm{d}}_{\mathscr{A}})$ is unique up to a unique isomorphism: given any pair $(\mathscr{M},{\mathrm{d}})$ with the same properties, there exists a unique isomorphism of $L^{1}(\mathfrak{m})$-Banach $L^{\infty}(\mathfrak{m})$-modules $\Phi\colon L^{1}(T^{*}{\rm X};\mathscr{A})\to\mathscr{M}$ such that ${\mathrm{d}}=\Phi\circ{\mathrm{d}}_{\mathscr{A}}$. Moreover, ${\mathrm{d}}_{\mathscr{A}}$ satisfies the Leibniz rule, i.e.
${\mathrm{d}}_{\mathscr{A}}(fg)=f\,{\mathrm{d}}_{\mathscr{A}}g+g\,{\mathrm{d}}_{\mathscr{A}}f\quad\text{ for every }f,g\in\mathscr{A}.$ (3.5)
Proof.
This kind of construction is due to Gigli [23]; see also [22] or [24]. Technically speaking, the stated existence and uniqueness are direct consequences of [29, Theorem 3.19]. The Leibniz rule (3.5) follows from the locality (3.4) of $|{\rm D}f|_{*,\mathscr{A}}$, arguing as in [23, Theorem 2.2.6]. ∎
In analogy with [23, Definition 2.3.1], we then define the Sobolev $\mathscr{A}$-tangent module as
$L^{\infty}(T{\rm X};\mathscr{A})\coloneqq L^{1}(T^{*}{\rm X};\mathscr{A})^{*}.$
Recall that $L^{\infty}(T{\rm X};\mathscr{A})$ is an $L^{\infty}(\mathfrak{m})$-Banach $L^{\infty}(\mathfrak{m})$-module. Recall also that the operator
$\textsc{Int}_{\mathscr{A}}\coloneqq\textsc{Int}_{L^{1}(T^{*}{\rm X};\mathscr{A})}\colon L^{\infty}(T{\rm X};\mathscr{A})\to L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}$
is an isometric isomorphism of Banach spaces, see (2.4). To be consistent with [23], we will denote by $\omega(v)\in L^{1}(\mathfrak{m})$ (instead of $v(\omega)\in L^{1}(\mathfrak{m})$) the ‘pointwise duality pairing’ between two elements $\omega\in L^{1}(T^{*}{\rm X};\mathscr{A})$ and $v\in L^{\infty}(T{\rm X};\mathscr{A})$, even though $L^{\infty}(T{\rm X};\mathscr{A})$ is the module dual of $L^{1}(T^{*}{\rm X};\mathscr{A})$ (and, in general, not the other way round).
3.3. The space ${\rm BV}_{\rm W}$
A notion of metric BV space in terms of an integration-by-parts formula involving derivations was introduced in [17, Definition 3.1]; see also [16]. We generalise it as follows:
Definition 3.10 (BV space via derivations).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. Then we define ${\rm BV}_{\rm W}({\rm X};\mathscr{A})$ as the space of all functions $f\in L^{1}(\mathfrak{m})$ for which there exists a bounded linear operator ${\bf L}_{f}\colon{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})\to\mathcal{M}({\rm X})$ such that
[…]
Remark 3.11.
Given any function $f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A})$, the operator ${\bf L}_{f}$ is uniquely determined. Indeed, if another operator $\tilde{\bf L}_{f}$ satisfies the same properties, then (3.6b) and (3.6a) imply that
$\int h\,{\mathrm{d}}({\bf L}_{f}(b)-\tilde{\bf L}_{f}(b))={\bf L}_{f}(hb)({\rm X})-\tilde{\bf L}_{f}(hb)({\rm X})=-\int f\,{\rm div}(hb)\,{\mathrm{d}}\mathfrak{m}+\int f\,{\rm div}(hb)\,{\mathrm{d}}\mathfrak{m}=0$
for every $h\in\mathscr{A}$ and $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$, whence it follows that ${\bf L}_{f}=\tilde{\bf L}_{f}$ thanks to (2.2). $\blacksquare$
By Remark 3.11, it makes sense to define the total $\mathscr{A}$-variation of any $f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A})$ as
${\rm V}_{\mathscr{A}}(f)\coloneqq\|{\bf L}_{f}\|_{\mathcal{L}(\overline{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A});\mathcal{M}({\rm X}))},$
where we keep the notation ${\bf L}_{f}$ to indicate the unique bounded linear operator from $\overline{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ to $\mathcal{M}({\rm X})$ that extends ${\bf L}_{f}$. One can readily check that ${\rm BV}_{\rm W}({\rm X};\mathscr{A})$ is a Banach space if endowed with the norm
$\|f\|_{{\rm BV}_{\rm W}({\rm X};\mathscr{A})}\coloneqq\|f\|_{L^{1}(\mathfrak{m})}+{\rm V}_{\mathscr{A}}(f)\quad\text{ for every }f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A}).$
Proposition 3.12.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. For any $f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A})$, we define its total $\mathscr{A}$-variation measure as
$|{\bf D}f|_{\mathscr{A}}\coloneqq\bigvee\big{\{}|{\bf L}_{f}(b)|\;\big{|}\;b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}),\,|b|_{\mathscr{A}}\leq 1\big{\}}.$
Then $|{\bf D}f|_{\mathscr{A}}$ is a finite non-negative Borel measure on ${\rm X}$ such that $|{\bf D}f|_{\mathscr{A}}({\rm X})={\rm V}_{\mathscr{A}}(f)$.
Remark 3.13.
Since ${\bf L}_{f}(0)$ is the null measure and ${\bf L}_{f}(-b)=-{\bf L}_{f}(b)$ for every $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$, the definition of $|{\bf D}f|_{\mathscr{A}}$ can be equivalently rewritten as
$|{\bf D}f|_{\mathscr{A}}(E)=\sup\sum_{i=1}^{n}{\bf L}_{f}(b_{i})(E_{i})\quad\text{ for every Borel set }E\subseteq{\rm X},$
where the supremum is taken among all finite Borel partitions $(E_{i})_{i=1}^{n}$ of $E$ and all collections $(b_{i})_{i=1}^{n}\subseteq{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ such that $|b_{i}|_{\mathscr{A}}\leq 1$ and $E_{i}$ is contained in the support of the positive part of ${\bf L}_{f}(b_{i})$. This is because, given $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ and a Borel subset $F\subseteq{\rm X}$, we have that
$|{\bf L}_{f}(b)|(F)={\bf L}_{f}(b)(F\cap P_{b})+{\bf L}_{f}(-b)(F\setminus P_{b}),$
where $P_{b}$ is the support of the positive part of ${\bf L}_{f}(b)$. Thus one can always replace the partition $(E_{i})_{i=1}^{n}$ with $(E_{i}\cap P_{b_{i}})_{i=1}^{n}\cup(E_{i}\setminus P_{b_{i}})_{i=1}^{n}$, and the collection $(b_{i})_{i=1}^{n}$ with $(b_{i})_{i=1}^{n}\cup(-b_{i})_{i=1}^{n}$. $\blacksquare$
Proof of Proposition 3.12.
Let $\lambda<|{\bf D}f|_{\mathscr{A}}({\rm X})$ be fixed. Taking into account Remark 3.13, we can find a Borel partition $E_{1},\ldots,E_{n}$ of ${\rm X}$ and $b_{1},\ldots,b_{n}\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ such that $|b_{i}|_{\mathscr{A}}\leq 1$ for all $i=1,\ldots,n$ and $\sum_{i=1}^{n}{\bf L}_{f}(b_{i})(E_{i})>\lambda$. Letting $\mu\coloneqq\sum_{i=1}^{n}|{\bf L}_{f}(b_{i})|$ and $\delta\coloneqq\sum_{i=1}^{n}{\bf L}_{f}(b_{i})(E_{i})-\lambda$, we know from Lemma 2.4 that there exist non-negative functions $\eta_{1},\ldots,\eta_{n}\in\mathscr{A}$ with $\sum_{i=1}^{n}\eta_{i}\leq 1$ such that $\|\eta_{i}-\mathbbm{1}_{E_{i}}\|_{L^{1}(\mu)}\leq\delta/n$ for every $i=1,\ldots,n$. Letting $b\coloneqq\sum_{i=1}^{n}\eta_{i}b_{i}\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$, we have that $|b|_{\mathscr{A}}\leq\sum_{i=1}^{n}\eta_{i}|b_{i}|_{\mathscr{A}}\leq 1$ and
$\bigg{|}\sum_{i=1}^{n}{\bf L}_{f}(b_{i})(E_{i})-{\bf L}_{f}(b)({\rm X})\bigg{|}\leq\sum_{i=1}^{n}\int|\mathbbm{1}_{E_{i}}-\eta_{i}|\,{\mathrm{d}}|{\bf L}_{f}(b_{i})|\leq\delta,$
whence it follows that
$\lambda=\sum_{i=1}^{n}{\bf L}_{f}(b_{i})(E_{i})-\delta\leq{\bf L}_{f}(b)({\rm X})\leq|{\bf L}_{f}(b)|({\rm X})\leq{\rm V}_{\mathscr{A}}(f).$
Letting $\lambda\nearrow|{\bf D}f|_{\mathscr{A}}({\rm X})$, we deduce that $|{\bf D}f|_{\mathscr{A}}({\rm X})\leq{\rm V}_{\mathscr{A}}(f)$, thus $|{\bf D}f|_{\mathscr{A}}$ is finite. Finally, we have $|{\bf L}_{f}(b)|({\rm X})\leq|{\bf D}f|_{\mathscr{A}}({\rm X})$ for all $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ with $|b|_{\mathscr{A}}\leq 1$, thus ${\rm V}_{\mathscr{A}}(f)\leq|{\bf D}f|_{\mathscr{A}}({\rm X})$. ∎
Remark 3.14.
Assuming in addition that $({\rm X},{\sf d})$ is complete, it holds that
$|{\bf D}f|_{\mathscr{A}}\text{ is a Radon measure for every }f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A}).$
Indeed, $|{\bf D}f|_{\mathscr{A}}$ is concentrated on ${\rm spt}(\mathfrak{m})$ as a consequence of (3.6c), thus it is a Radon measure because ${\rm spt}(\mathfrak{m})$ is complete and separable. $\blacksquare$
When $\mathscr{A}={\rm LIP}_{\star}({\rm X})$, we just write ${\rm BV}_{\rm W}({\rm X})$ and $|{\bf D}f|$ for brevity.
Lemma 3.15.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Assume that $({\rm X},{\sf d})$ is either complete or a Radon space. Then it holds that ${\rm BV}_{\rm H}({\rm X})\subseteq{\rm BV}_{\rm W}({\rm X})$ and
$|{\bf D}f|\leq|{\bf D}f|_{*}\quad\text{ for every }f\in{\rm BV}_{\rm H}({\rm X}).$
Proof.
Let $f\in{\rm BV}_{\rm H}({\rm X})$ be given. Fix any $b\in{\rm Der}^{\infty}_{\infty}({\rm X})$. Take a sequence $(f_{n})_{n}\subseteq{\rm LIP}_{\star}({\rm X})$ such that $f_{n}\to f$ in $L^{1}(\mathfrak{m})$ and $\int{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}\to\|{\bf D}f\|_{*}$. Accounting for Remark 3.3, we have that ${\rm lip}_{a}(f_{n})\mathfrak{m}\rightharpoonup|{\bf D}f|_{*}$ weakly. Since $|{\bf D}f|_{*}$ is a Radon measure by Remark 3.4, we deduce from [9, Theorem 8.6.4] that $({\rm lip}_{a}(f_{n})\mathfrak{m})_{n}$ is a tight sequence. As $|b(f_{n})\mathfrak{m}|=|b(f_{n})|\mathfrak{m}\leq|b|{\rm lip}_{a}(f_{n})\mathfrak{m}$ for every $n\in\mathbb{N}$, we deduce that $(b(f_{n})\mathfrak{m})_{n}$ is a tight sequence as well. Since $(b(f_{n})\mathfrak{m})_{n}$ is also bounded in total variation norm, by [9, Theorem 8.6.7] we deduce that, up to a non-relabelled subsequence, $b(f_{n})\mathfrak{m}\rightharpoonup{\bf L}_{f}(b)$ weakly for some ${\bf L}_{f}(b)\in\mathcal{M}({\rm X})$. For any $h\in{\rm LIP}_{\star}({\rm X})$, we can compute
$\int h\,{\mathrm{d}}{\bf L}_{f}(b)=\lim_{n\to\infty}\int h\,b(f_{n})\,{\mathrm{d}}\mathfrak{m}=-\lim_{n\to\infty}\int f_{n}(h\,{\rm div}(b)+b(h))\,{\mathrm{d}}\mathfrak{m}=-\int f(h\,{\rm div}(b)+b(h))\,{\mathrm{d}}\mathfrak{m}.$
Therefore, the measure ${\bf L}_{f}(b)$ is independent of the approximating sequence and $b(f_{n})\mathfrak{m}\rightharpoonup{\bf L}_{f}(b)$ holds for the original sequence $(f_{n})_{n}$. In particular, we have that ${\bf L}_{f}(b)$ is concentrated on ${\rm spt}(\mathfrak{m})$, proving (3.6c). If $b,\tilde{b}\in{\rm Der}^{\infty}_{\infty}({\rm X})$ and $\lambda\in\mathbb{R}$, then $(\lambda b+\tilde{b})(f_{n})\mathfrak{m}=\lambda\,b(f_{n})\mathfrak{m}+\tilde{b}(f_{n})\mathfrak{m}$ weakly converges both to ${\bf L}_{f}(\lambda b+\tilde{b})$ and $\lambda\,{\bf L}_{f}(b)+{\bf L}_{f}(\tilde{b})$, so that ${\bf L}_{f}$ is linear. Given $h,\tilde{h}\in{\rm LIP}_{\star}({\rm X})$, we have that $\int\tilde{h}\,{\mathrm{d}}{\bf L}_{f}(hb)=\lim_{n}\int\tilde{h}(hb)(f_{n})\,{\mathrm{d}}\mathfrak{m}=\lim_{n}\int(\tilde{h}h)b(f_{n})\,{\mathrm{d}}\mathfrak{m}=\int\tilde{h}h\,{\mathrm{d}}{\bf L}_{f}(b)=\int\tilde{h}\,{\mathrm{d}}(h\,{\bf L}_{f}(b))$, which yields (3.6b) thanks to (2.2). Moreover, for every $b\in{\rm Der}^{\infty}_{\infty}({\rm X})$ we can compute
${\bf L}_{f}(b)({\rm X})=\lim_{n\to\infty}\int b(f_{n})\,{\mathrm{d}}\mathfrak{m}=-\lim_{n\to\infty}\int f_{n}\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m}=-\int f\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m},$
which gives (3.6a). Next, fix a derivation $b\in{\rm Der}^{\infty}_{\infty}({\rm X})$ with $|b|\leq 1$ and a compact set $K\subseteq{\rm X}$. We claim that
${\bf L}_{f}(b)(K)\leq|{\bf D}f|_{*}(K),$ (3.7)
whence (since $|{\bf L}_{f}(b)|$ is a Radon measure) it follows that $f\in{\rm BV}_{\rm W}({\rm X})$ and $|{\bf D}f|\leq|{\bf D}f|_{*}$. To prove the claim, pick a sequence of open sets $(U_{i})_{i}$ such that $K\subset U_{i}$ for every $i\in\mathbb{N}$ and satisfying ${\bf L}_{f}(b)(U_{i})\to{\bf L}_{f}(b)(K)$, $|{\bf D}f|_{*}(U_{i})\to|{\bf D}f|_{*}(K)$. Now fix $i\in\mathbb{N}$ and take a non-decreasing sequence $(\eta_{i,j})_{j}$ of Lipschitz cut-off functions $\eta_{i,j}\colon{\rm X}\to[0,1]$ such that $\eta_{i,j}(x)\nearrow\mathbbm{1}_{U_{i}}(x)$ as $j\to\infty$ for every $x\in{\rm X}$. Note that
$\begin{split}\int\eta_{i,j}\,{\mathrm{d}}{\bf L}_{f}(b)&=\lim_{n\to\infty}\int\eta_{i,j}\,b(f_{n})\,{\mathrm{d}}\mathfrak{m}\leq\liminf_{n\to\infty}\int\eta_{i,j}|b|\,{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}\\ &\leq\liminf_{n\to\infty}\int\eta_{i,j}\,{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}=\int\eta_{i,j}\,{\mathrm{d}}|{\bf D}f|_{*}\leq|{\bf D}f|_{*}(U_{i})\end{split}$
for every $i,j\in\mathbb{N}$. Applying the dominated convergence theorem, we deduce that
${\bf L}_{f}(b)(U_{i})=\lim_{j\to\infty}\int\eta_{i,j}\,{\mathrm{d}}{\bf L}_{f}(b)\leq|{\bf D}f|_{*}(U_{i})\quad\text{ for every }i\in\mathbb{N}.$
Therefore, we can finally conclude that
${\bf L}_{f}(b)(K)=\lim_{i\to\infty}{\bf L}_{f}(b)(U_{i})\leq\lim_{i\to\infty}|{\bf D}f|_{*}(U_{i})=|{\bf D}f|_{*}(K),$
which gives (3.7). Thus, the proof is complete. ∎
Remark 3.16.
In the case where $({\rm X},{\sf d})$ is complete, by slightly adapting the proof of Lemma 3.15 (and taking also Proposition 3.5 into account) one can prove that ${\rm BV}_{\rm H}({\rm X})\subseteq{\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X}))$ and $|{\bf D}f|_{{\rm LIP}_{b}({\rm X})}\leq|{\bf D}f|_{*}$ for every $f\in{\rm BV}_{\rm H}({\rm X})$. $\blacksquare$
4. Equivalence results
Given a metric space $({\rm X},{\sf d})$ and $x\in{\rm X}$, we define the $1$-Lipschitz function ${\sf d}_{x}\colon{\rm X}\to[0,+\infty)$ as
${\sf d}_{x}(y)\coloneqq{\sf d}(x,y)\quad\text{ for every }y\in{\rm X}.$
Definition 4.1 (Good algebra).
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{b}({\rm X})$. Then we say that $\mathscr{A}$ is a good algebra if, given any $x\in{\rm X}$ and $r>0$, there exist $(f_{n})_{n}\subseteq\mathscr{A}$ and $\rho\in L^{\infty}(\mathfrak{m})$ with $0\leq\rho\leq 1$ such that $\|f_{n}-{\sf d}_{x}\wedge r\|_{L^{1}(\mathfrak{m})}\to 0$ and ${\rm lip}_{a}(f_{n})\rightharpoonup\rho$ weakly in $L^{1}(\mathfrak{m})$.
Note that a unital separating subalgebra $\mathscr{A}$ of ${\rm LIP}_{b}({\rm X})$ is a good algebra precisely when
${\sf d}_{x}\wedge r\in{\rm H}^{1,1}({\rm X};\mathscr{A}),\qquad|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}\leq 1\quad\text{ for every }x\in{\rm X}\text{ and }r>0.$
Clearly, ${\rm LIP}_{b}({\rm X})$ itself is a good algebra.
Example 4.2 (Some good algebras).
In [36] is introduced a notion of compatible algebra in the setting of extended metric-topological measure spaces. These are Hausdorff topological spaces $({\rm X},\tau)$ endowed with a (possibly extended) metric ${\sf d}$ which has good compatibility properties with $\tau$, together with a finite, non-negative, Radon measure $\mathfrak{m}$ (on $({\rm X},\tau)$). Whenever $\tau$ coincides with the topology induced by ${\sf d}$ and ${\sf d}$ is finite, we recover our setting, i.e. $({\rm X},{\sf d},\mathfrak{m})$ is a metric measure space in the sense of Definition 2.1. In this case, an algebra $\mathscr{A}$ is said to be compatible if
${\sf d}(x,y)=\sup\left\{\frac{|f(x)-f(y)|}{{\rm Lip}(f)}\;\bigg{|}\;f\in\mathscr{A},\,{\rm Lip}(f)\neq 0\right\}\quad\text{ for every }x,y\in{\rm X}.$
It is not difficult to check that this implies that the functions ${\sf d}_{x}$, $x\in{\rm X}$ can be approximated in the following sense: there exists a sequence $(f_{n})_{n}\subset\mathscr{A}$ such that
$f_{n}\to{\sf d}_{x}\quad\text{$\mathfrak{m}$-a.e.},\quad{\rm lip}_{a}(f_{n})\rightharpoonup\rho\text{ in }L^{1}(\mathfrak{m})\quad\text{ for some }0\leq\rho\leq 1.$
This is precisely (even if there is considered only for $p>1$) the density condition in [20, Theorem 2.13]. In turn this last condition implies that $\mathscr{A}$ is a good algebra according to Definition 4.1, simply using a truncation argument of the distance. Therefore, a variety of algebras in many metric spaces are good algebras:
  1. (1)
    Smooth functions in Riemannian manifolds and Euclidean spaces [20, Remark 2.21].
  2. (2)
    Cylinder functions in Banach and Hilbert spaces [36, Example 2.1.19].
  3. (3)
    Algebras generated by approximating distances [36, Example 2.1.20].
  4. (4)
    Cylinder functions in the Wasserstein, Hellinger and Hellinger–Kantorovich metric spaces of (probability) measures [37, 20, 15]. $\blacksquare$
Remark 4.3.
If $\mathscr{A}$ is a good algebra, then the approximation $(f_{n})_{n}$ of ${\sf d}_{x}\wedge r$ can be also chosen so that $0\leq f_{n}\leq r$ for every $n\in\mathbb{N}$. Indeed, letting $(f_{n})_{n}$ be as in Definition 4.1, we have that $u_{n}\coloneqq(f_{n}\wedge r)\vee 0\to{\sf d}_{x}\wedge r$ in $L^{1}(\mathfrak{m})$. Thanks to [36, Lemma 2.1.26], we can find a sequence $(\tilde{g}_{n})_{n}\subseteq\mathscr{A}$ such that $|\tilde{g}_{n}-u_{n}|\leq\frac{1}{n}$ and ${\rm lip}_{a}(\tilde{g}_{n})\leq{\rm lip}_{a}(u_{n})$ for all $n\in\mathbb{N}$. By the Dunford–Pettis theorem, we thus have (up to a non-relabelled subsequence) that ${\rm lip}_{a}(\tilde{g}_{n})\rightharpoonup\rho$ weakly in $L^{1}(\mathfrak{m})$, for some function $\rho\in L^{\infty}(\mathfrak{m})$ with $0\leq\rho\leq 1$. Finally, letting $g_{n}\coloneqq r\big{(}r+\frac{2}{n}\big{)}^{-1}\big{(}\tilde{g}_{n}+\frac{1}{n}\big{)}$ for every $n\in\mathbb{N}$, we obtain a sequence $(g_{n})_{n}\subseteq\mathscr{A}$ with $0\leq g_{n}\leq r$ such that $g_{n}\to{\sf d}_{x}\wedge r$ strongly in $L^{1}(\mathfrak{m})$ and ${\rm lip}_{a}(g_{n})\rightharpoonup\rho$ weakly in $L^{1}(\mathfrak{m})$, as desired. $\blacksquare$
Proposition 4.4.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space and $\mathscr{A}$ a good subalgebra of ${\rm LIP}_{b}({\rm X})$. Let $f\in{\rm LIP}_{b}({\rm X})$ be given. Then there exists a sequence $(f_{n})_{n}\subseteq\mathscr{A}$ such that the following hold:
  • $\rm i)$
    $\inf_{X}f\leq f_{n}\leq\sup_{\rm X}f$ for every $n\in\mathbb{N}$.
  • $\rm ii)$
    $f_{n}\to f$ in $L^{1}(\mathfrak{m})$ as $n\to\infty$.
  • $\rm iii)$
    $({\rm lip}_{a}(f_{n}))_{n}$ is dominated in $L^{1}(\mathfrak{m})$, i.e. there exists $G\in L^{1}(\mathfrak{m})^{+}$ such that ${\rm lip}_{a}(f_{n})\leq G$ holds $\mathfrak{m}$-a.e. on ${\rm X}$ for every $n\in\mathbb{N}$.
  • $\rm iv)$
    $({\rm lip}_{a}(f_{n})-{\rm Lip}(f))^{+}\to 0$ in $L^{1}(\mathfrak{m})$ as $n\to\infty$.
Proof.
Without loss of generality, we can assume that ${\rm Lip}(f)>0$. Fix a sequence $(x_{k})_{k}\subseteq{\rm X}$ that is dense in ${\rm spt}(\mathfrak{m})$. Define $L\coloneqq{\rm Lip}(f)$, $m\coloneqq\inf_{\rm X}f$, $M\coloneqq\sup_{\rm X}f$ and
$\tilde{g}_{k}\coloneqq(f(x_{k})-L{\sf d}_{x_{k}})\vee m=f(x_{k})-L\bigg{(}{\sf d}_{x_{k}}\wedge\frac{f(x_{k})-m}{L}\bigg{)}\quad\text{ for every }k\in\mathbb{N}.$
Notice that $\tilde{g}_{k}\in{\rm LIP}_{b}({\rm X})$ and ${\rm Lip}(\tilde{g}_{k})\leq L$. Since $\mathscr{A}$ is a good algebra, we can find $(f^{j}_{k})_{j}\subseteq\mathscr{A}$ such that $f^{j}_{k}\to\tilde{g}_{k}$ in $L^{1}(\mathfrak{m})$ and ${\rm lip}_{a}(f^{j}_{k})\rightharpoonup|{\rm D}\tilde{g}_{k}|_{*,\mathscr{A}}\leq L$ in $L^{1}(\mathfrak{m})$ as $j\to\infty$. By Remark 4.3, we can assume that $m\leq f^{j}_{k}\leq M$ for every $k,j\in\mathbb{N}$. Define
$g_{n}\coloneqq\tilde{g}_{1}\vee\ldots\vee\tilde{g}_{n}\in{\rm LIP}_{b}({\rm X})\quad\text{ for every }n\in\mathbb{N}.$
Given that $m\leq g_{n}\leq M$ for every $n\in\mathbb{N}$ and $g_{n}(x)\to f(x)$ for every $x\in{\rm spt}(\mathfrak{m})$, we have that $g_{n}\to f$ in $L^{1}(\mathfrak{m})$ by the dominated convergence theorem. Now fix $n\in\mathbb{N}$. Since $f^{j}_{1}\vee\ldots\vee f^{j}_{n}\to g_{n}$ in $L^{1}(\mathfrak{m})$ as $j\to\infty$, there exists $j(n)\in\mathbb{N}$ such that $\|f^{j(n)}_{1}\vee\ldots\vee f^{j(n)}_{n}-g_{n}\|_{L^{1}(\mathfrak{m})}\leq 1/n$ and
$\|h_{n}^{k}\|_{L^{1}(\mathfrak{m})}\leq\frac{1}{n2^{n}}\quad\text{ for every }k=1,\ldots,n,\text{ where }h_{n}^{k}\coloneqq\big{|}{\rm lip}_{a}(f_{k}^{j(n)})-|{\rm D}\tilde{g}_{k}|_{*,\mathscr{A}}\big{|}.$
By [36, Lemma 2.1.26], we can find a function $\tilde{f}_{n}\in\mathscr{A}$ such that $\sup_{\rm X}|f^{j(n)}_{1}\vee\ldots\vee f^{j(n)}_{n}-\tilde{f}_{n}|\leq 1/n$ and ${\rm lip}_{a}(\tilde{f}_{n})\leq{\rm lip}_{a}(f_{1}^{j(n)})\vee\ldots\vee{\rm lip}_{a}(f_{n}^{j(n)})$. In particular, $m-\frac{1}{n}\leq\tilde{f}_{n}\leq M+\frac{1}{n}$ for every $n\in\mathbb{N}$. It also follows that $\|\tilde{f}_{n}-g_{n}\|_{L^{1}(\mathfrak{m})}\leq(1+\mathfrak{m}({\rm X}))/n$ and
${\rm lip}_{a}(\tilde{f}_{n})\leq(|{\rm D}\tilde{g}_{1}|_{*,\mathscr{A}}+h_{n}^{1})\vee\dots\vee(|{\rm D}\tilde{g}_{n}|_{*,\mathscr{A}}+h_{n}^{n})\leq L+h_{n}^{1}\vee\dots\vee h_{n}^{n}\leq L+h_{n}^{1}+\dots+h_{n}^{n}.$
Note that $h_{n}\coloneqq h_{n}^{1}+\dots+h_{n}^{n}$ satisfies $\|h_{n}\|_{L^{1}(\mathfrak{m})}\leq 2^{-n}$, thus $h\coloneqq\sum_{n=1}^{\infty}h_{n}$ belongs to $L^{1}(\mathfrak{m})$. Therefore, $\tilde{f}_{n}\to f$ in $L^{1}(\mathfrak{m})$ as $n\to\infty$ and ${\rm lip}_{a}(\tilde{f}_{n})\leq L+h\in L^{1}(\mathfrak{m})$ for every $n\in\mathbb{N}$. Moreover, we have that $({\rm lip}_{a}(\tilde{f}_{n})-L)^{+}\leq h_{n}\to 0$ in $L^{1}(\mathfrak{m})$ as $n\to\infty$. Finally, defining $(f_{n})_{n}\subseteq\mathscr{A}$ as
$f_{n}\coloneqq c_{n}\bigg{(}\tilde{f}_{n}-m+\frac{1}{n}\bigg{)}+m,\quad\text{ where we set }c_{n}\coloneqq\frac{M-m}{M-m+\frac{2}{n}},$
one can then easily check that $(f_{n})_{n}$ satisfies i), ii), iii) and iv). The statement is achieved. ∎
Theorem 4.5.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a good subalgebra of ${\rm LIP}_{b}({\rm X})$. Define the operator $\phi_{\mathscr{A}}\colon{\rm Der}^{\infty}_{\infty}({\rm X};{\rm LIP}_{b}({\rm X}))\to{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ as
$\phi_{\mathscr{A}}(b)\coloneqq b|_{\mathscr{A}}\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})\quad\text{ for every }b\in{\rm Der}^{\infty}_{\infty}({\rm X};{\rm LIP}_{b}({\rm X})).$
Then $\phi_{\mathscr{A}}$ is an isomorphism of $\mathscr{A}$-modules such that the identity ${\rm div}(\phi_{\mathscr{A}}(b))={\rm div}(b)$ holds for every $b\in{\rm Der}^{\infty}_{\infty}({\rm X};{\rm LIP}_{b}({\rm X}))$. Moreover, it holds that
$|\phi_{\mathscr{A}}(b)|_{\mathscr{A}}=|b|\quad\text{ for every }b\in{\rm Der}^{\infty}_{\infty}({\rm X};{\rm LIP}_{b}({\rm X})).$
Proof.
Let us write ${\rm L}_{\rm X}\coloneqq{\rm LIP}_{b}({\rm X})$ for brevity. First of all, notice that $\phi_{\mathscr{A}}(b)\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ and $|\phi_{\mathscr{A}}(b)|_{\mathscr{A}}\leq|b|$ for every $b\in{\rm Der}^{\infty}_{\infty}({\rm X};{\rm L}_{\rm X})$. Moreover, the operator $\phi_{\mathscr{A}}$ is a homomorphism of $\mathscr{A}$-modules by construction.
To conclude, it remains to show that if $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ is given, there exists $\bar{b}\in{\rm Der}^{\infty}_{\infty}({\rm X};{\rm L}_{\rm X})$ such that $\phi_{\mathscr{A}}(\bar{b})=b$ and $|\bar{b}|\leq|b|_{\mathscr{A}}$. To prove it, fix a function $f\in{\rm L}_{\rm X}$ and take a sequence $(f_{n})_{n}\subseteq\mathscr{A}$ that approximates $f$ in the sense of Proposition 4.4. Since $|b(f_{n})|\leq|b|_{\mathscr{A}}{\rm lip}_{a}(f_{n})$ for every $n\in\mathbb{N}$, we have that $(b(f_{n}))_{n}$ is dominated in $L^{1}(\mathfrak{m})$, thus the Dunford–Pettis theorem ensures that we can extract a subsequence $(f_{n_{k}})_{k}$ such that $b(f_{n_{k}})\rightharpoonup\bar{b}(f)$ in $L^{1}(\mathfrak{m})$, for some limit function $\bar{b}(f)\in L^{1}(\mathfrak{m})$. Note that whenever a sequence $(\tilde{f}_{i})_{i}\subseteq\mathscr{A}$ and a function $h\in L^{1}(\mathfrak{m})$ satisfy $\tilde{f}_{i}\rightharpoonup f$ and $b(\tilde{f}_{i})\rightharpoonup h$ weakly in $L^{1}(\mathfrak{m})$, for any $g\in\mathscr{A}$ we have that
$\begin{split}\int hg\,{\mathrm{d}}\mathfrak{m}&=\lim_{i\to\infty}\int g\,b(\tilde{f}_{i})\,{\mathrm{d}}\mathfrak{m}=\lim_{i\to\infty}\int b(g\tilde{f}_{i})-\tilde{f}_{i}\,b(g)\,{\mathrm{d}}\mathfrak{m}\\ &=-\lim_{i\to\infty}\int\tilde{f}_{i}(g\,{\rm div}(b)+b(g))\,{\mathrm{d}}\mathfrak{m}=-\int f(g\,{\rm div}(b)+b(g))\,{\mathrm{d}}\mathfrak{m}.\end{split}$ (4.1)
This shows that $\bar{b}(f)$ is independent of the chosen subsequence $(f_{n_{k}})_{k}$ and that the following holds: if $(\tilde{f}_{i})_{i}\subseteq\mathscr{A}$ satisfies $\tilde{f}_{i}\rightharpoonup f$ weakly in $L^{1}(\mathfrak{m})$ and $(b(\tilde{f}_{i}))_{i}$ is dominated in $L^{1}(\mathfrak{m})$, then $b(\tilde{f}_{i})\rightharpoonup\bar{b}(f)$ weakly in $L^{1}(\mathfrak{m})$. In particular, we have that $\bar{b}|_{\mathscr{A}}=b$. Let us now check that $\bar{b}\in{\rm Der}^{\infty}_{\infty}({\rm X};{\rm L}_{\rm X})$:
  • Fix $f,g\in{\rm L}_{\rm X}$ and $\alpha,\beta\in\mathbb{R}$. Let $(f_{n})_{n}\subseteq\mathscr{A}$ (resp. $(g_{n})_{n}\subseteq\mathscr{A}$) be an approximating sequence for $f$ (resp. $g$) as in Proposition 4.4. Letting $n\to\infty$ in the identity $b(\alpha f_{n}+\beta g_{n})=\alpha\,b(f_{n})+\beta\,b(g_{n})$, we obtain that $\bar{b}(\alpha f+\beta g)=\alpha\,\bar{b}(f)+\beta\,\bar{b}(g)$, proving that $\bar{b}$ is linear.
  • Note that whenever $f\in{\rm L}_{\rm X}$ and $(f_{n})_{n}\subseteq\mathscr{A}$ is an approximating sequence for $f$ as in Proposition 4.4, we have that
    $\int\bar{b}(f)\,{\mathrm{d}}\mathfrak{m}=\lim_{n\to\infty}\int b(f_{n})\,{\mathrm{d}}\mathfrak{m}=-\lim_{n\to\infty}\int f_{n}\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m}=-\int f\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m}.$ (4.2)
  • We claim that
    $\bar{b}(fg)=f\,b(g)+g\,\bar{b}(f)\quad\text{ for every }f\in{\rm L}_{\rm X}\text{ and }g\in\mathscr{A}.$ (4.3)
    Indeed, letting $(f_{n})_{n}\subseteq\mathscr{A}$ be an approximating sequence for $f$ as in Proposition 4.4, we have that $b(f_{n}g)=f_{n}\,b(g)+g\,b(f_{n})$ for every $n\in\mathbb{N}$, which gives $\bar{b}(fg)=f\,b(g)+g\,\bar{b}(f)$.
  • Next, we claim that $\bar{b}$ has the following property: if $f\in{\rm L}_{\rm X}$ and $(\tilde{f}_{i})_{i}\subseteq{\rm L}_{\rm X}$ are such that $\tilde{f}_{i}\rightharpoonup f$ weakly in $L^{1}(\mathfrak{m})$ and $(\bar{b}(\tilde{f}_{i}))_{i}$ is dominated in $L^{1}(\mathfrak{m})$, then $\bar{b}(\tilde{f}_{i})\rightharpoonup\bar{b}(f)$ weakly in $L^{1}(\mathfrak{m})$. To prove it, one can argue as in (4.1), by using (4.2) and (4.3).
  • Consequently, we can show that $\bar{b}$ satisfies the Leibniz rule, i.e.
    $\bar{b}(fg)=f\,\bar{b}(g)+g\,\bar{b}(f)\quad\text{ for every }f,g\in{\rm L}_{\rm X}.$ (4.4)
    Indeed, letting $(g_{n})_{n}\subseteq\mathscr{A}$ be an approximating sequence for $g$ as in Proposition 4.4, we know from (4.3) that $\bar{b}(fg_{n})=f\,b(g_{n})+g_{n}\bar{b}(f)$ for every $n\in\mathbb{N}$, whence (4.4) follows.
  • We also claim that
    $|\bar{b}(f)|\leq{\rm Lip}(f)|b|_{\mathscr{A}}\quad\text{ for every }f\in{\rm L}_{\rm X}.$ (4.5)
    To prove it, take an approximating sequence $(f_{n})_{n}\subseteq\mathscr{A}$ for $f$ as in Proposition 4.4 and set $h_{n}\coloneqq({\rm lip}_{a}(f_{n})-{\rm Lip}(f))^{+}$. Letting $n\to\infty$ in $\bar{b}(f_{n})\leq|b|_{\mathscr{A}}{\rm lip}_{a}(f_{n})\leq|b|_{\mathscr{A}}({\rm Lip}(f)+h_{n})$, we get $\bar{b}(f)\leq{\rm Lip}(f)|b|_{\mathscr{A}}$. Doing the same for $-f$ instead, we obtain that $-\bar{b}(f)\leq{\rm Lip}(f)|b|_{\mathscr{A}}$, proving (4.5).
  • Fix an equi-bounded, equi-Lipschitz sequence $(f_{n})_{n}\subseteq{\rm L}_{\rm X}$ converging pointwise to $f\in{\rm L}_{\rm X}$. Since $|\bar{b}(f_{n})|\leq{\rm Lip}(f_{n})|b|_{\mathscr{A}}$ by (4.5), the sequence $(\bar{b}(f_{n}))_{n}$ is bounded in $L^{\infty}(\mathfrak{m})$, thus it is dominated in $L^{1}(\mathfrak{m})$. Given that $f_{n}\to f$ in $L^{1}(\mathfrak{m})$ (by the dominated convergence theorem), we have that $\bar{b}(f_{n})\rightharpoonup\bar{b}(f)$ weakly in $L^{1}(\mathfrak{m})$ by the fourth point above. Since $L^{1}(\mathfrak{m})$ is separable, by the Banach–Alaoglu theorem we can extract a subsequence $(f_{n_{k}})_{k}$ such that $(\bar{b}(f_{n_{k}}))_{k}$ has a weak limit $G$ in $L^{\infty}(\mathfrak{m})$. In particular, $\bar{b}(f_{n_{k}})\rightharpoonup G$ weakly in $L^{1}(\mathfrak{m})$, thus $G=\bar{b}(f)$ and accordingly the original sequence $\bar{b}(f_{n})$ weakly converges to $\bar{b}(f)$ in $L^{\infty}(\mathfrak{m})$. This proves that $\bar{b}$ is weakly continuous (in the sense of [7, Lemma 4.12 i)]). By suitably adapting the proof of [7, Corollary 4.14] (exploiting the fact that ${\rm spt}(\mathfrak{m})$ is separable), we can conclude that $|\bar{b}(f)|\leq|b|_{\mathscr{A}}{\rm lip}_{a}(f)$ for every $f\in{\rm L}_{\rm X}$.
All in all, we showed that $\bar{b}\in{\rm Der}^{\infty}_{\infty}({\rm X};{\rm L}_{\rm X})$, $\phi_{\mathscr{A}}(\bar{b})=b$ and $|\bar{b}|\leq|b|_{\mathscr{A}}$. The proof is complete. ∎
Corollary 4.6.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a good subalgebra of ${\rm LIP}_{b}({\rm X})$. Then ${\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X}))\subseteq{\rm BV}_{\rm W}({\rm X};\mathscr{A})$ and $|{\bf D}f|_{\mathscr{A}}\leq|{\bf D}f|_{{\rm LIP}_{b}({\rm X})}$ for all $f\in{\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X}))$.
Proof.
Let $f\in{\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X}))$ and let ${\bf L}_{f}\colon{\rm Der}^{\infty}_{\infty}({\rm X};{\rm LIP}_{b}({\rm X}))\to\mathcal{M}({\rm X})$ be as in Definition 3.10 for $f$ and the algebra ${\rm LIP}_{b}({\rm X})$. We define the operator $\tilde{\bf L}_{f}\colon{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})\to\mathcal{M}({\rm X})$ as
$\tilde{\bf L}_{f}(b):={\bf L}_{f}(\phi_{\mathscr{A}}^{-1}(b))\quad\text{ for every }b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}),$
where $\phi_{\mathscr{A}}$ is as in Theorem 4.5. It is easy to check that $\tilde{\bf L}_{f}$ satisfies Definition 3.10 for $f$ and the algebra $\mathscr{A}$. The inequality between the total variation measures follows immediately from the properties of $\phi_{\mathscr{A}}$. ∎
Let us now fix a unital separating subalgebra $\mathscr{A}$ of ${\rm LIP}_{\star}({\rm X})$. We can regard ${\mathrm{d}}_{\mathscr{A}}$ as an unbounded linear operator ${\mathrm{d}}_{\mathscr{A}}\colon L^{1}(\mathfrak{m})\to L^{1}(T^{*}{\rm X};\mathscr{A})$ whose domain is $D({\mathrm{d}}_{\mathscr{A}})=\mathscr{A}$. Given that $\mathscr{A}$ is dense in $L^{1}(\mathfrak{m})$, we have that ${\mathrm{d}}_{\mathscr{A}}$ is densely defined, thus accordingly its adjoint operator
${\mathrm{d}}_{\mathscr{A}}^{*}\colon L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}\to L^{\infty}(\mathfrak{m})$
is well defined. Recall that the domain $D({\mathrm{d}}_{\mathscr{A}}^{*})$ of ${\mathrm{d}}_{\mathscr{A}}^{*}$ is a vector subspace of $L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}$, and that ${\mathrm{d}}_{\mathscr{A}}^{*}$ is characterised by $\int f\,{\mathrm{d}}_{\mathscr{A}}^{*}V\,{\mathrm{d}}\mathfrak{m}=V({\mathrm{d}}_{\mathscr{A}}f)$ for every $V\in D({\mathrm{d}}_{\mathscr{A}}^{*})$ and $f\in\mathscr{A}$.
Lemma 4.7.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. Fix any $v\in L^{\infty}(T{\rm X};\mathscr{A})$ with $\textsc{Int}_{\mathscr{A}}(v)\in D({\mathrm{d}}_{\mathscr{A}}^{*})$. Let us define $b_{v}\colon\mathscr{A}\to L^{1}(\mathfrak{m})$ as
$b_{v}(f)\coloneqq{\mathrm{d}}_{\mathscr{A}}f(v)\in L^{1}(\mathfrak{m})\quad\text{ for every }f\in\mathscr{A}.$
Then $b_{v}\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$. Moreover, it holds that $|b_{v}|_{\mathscr{A}}\leq|v|$ and ${\rm div}(b_{v})=-{\mathrm{d}}_{\mathscr{A}}^{*}(\textsc{Int}_{\mathscr{A}}(v))$.
Proof.
First, observe that ${\mathrm{d}}_{\mathscr{A}}f(v)\leq|v||{\mathrm{d}}_{\mathscr{A}}f|\leq|v|{\rm lip}_{a}(f)$ for every $f\in\mathscr{A}$, so that $b_{v}(f)\in L^{1}(\mathfrak{m})$ and $b_{v}$ satisfies Definition 2.6 i) with $G=|v|$. The linearity of $b_{v}$ follows from the linearity of ${\mathrm{d}}_{\mathscr{A}}$. For any $f,g\in\mathscr{A}$, we deduce from (3.5) that
$b_{v}(fg)={\mathrm{d}}_{\mathscr{A}}(fg)(v)=f\,{\mathrm{d}}_{\mathscr{A}}g(v)+g\,{\mathrm{d}}_{\mathscr{A}}f(v)=f\,b_{v}(g)+g\,b_{v}(f),$
which proves that $b_{v}$ satisfies Definition 2.6 ii), thus accordingly $b_{v}\in{\rm Der}^{\infty}({\rm X};\mathscr{A})$ and $|b_{v}|_{\mathscr{A}}\leq|v|$. Finally, recalling the definitions of ${\mathrm{d}}_{\mathscr{A}}^{*}$ and $\textsc{Int}_{\mathscr{A}}$, for any $f\in\mathscr{A}$ we can compute
$\int b_{v}(f)\,{\mathrm{d}}\mathfrak{m}=\int{\mathrm{d}}_{\mathscr{A}}f(v)\,{\mathrm{d}}\mathfrak{m}=\textsc{Int}_{\mathscr{A}}(v)({\mathrm{d}}_{\mathscr{A}}f)=\int f\,{\mathrm{d}}_{\mathscr{A}}^{*}(\textsc{Int}_{\mathscr{A}}(v))\,{\mathrm{d}}\mathfrak{m},$
which gives that $b_{v}\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ and ${\rm div}(b_{v})=-{\mathrm{d}}_{\mathscr{A}}^{*}(\textsc{Int}_{\mathscr{A}}(v))$. The proof is thus complete. ∎
We now define the pre-Cheeger $\mathscr{A}$-energy functional ~$\tilde{\mathcal{E}}_{\mathscr{A}}\colon L^{1}(\mathfrak{m})\to[0,+\infty]$ as
~$\tilde{\mathcal{E}}_{\mathscr{A}}(f)\coloneqq\left\{\begin{array}[]{ll}\int|{\rm D}f|_{*,\mathscr{A}}\,{\mathrm{d}}\mathfrak{m}\\ +\infty\end{array}\quad\begin{array}[]{ll}\text{ if }f\in\mathscr{A},\\ \text{ otherwise.}\end{array}\right.$
It can happen that ~$\tilde{\mathcal{E}}_{\mathscr{A}}|_{\mathscr{A}}\neq\mathcal{E}_{\mathscr{A}}|_{\mathscr{A}}$, but $\mathcal{E}_{\mathscr{A}}$ is the $L^{1}(\mathfrak{m})$-lower semicontinuous envelope of ~$\tilde{\mathcal{E}}_{\mathscr{A}}$ (thanks to (3.3) and to Remark 3.8). Notice also that the functional ~$\tilde{\mathcal{E}}_{\mathscr{A}}$ can be expressed as
~$\tilde{\mathcal{E}}_{\mathscr{A}}=\|\cdot\|_{L^{1}(T^{*}{\rm X};\mathscr{A})}\circ{\mathrm{d}}_{\mathscr{A}}.$ (4.6)
We denote by ${\sf n}_{\mathscr{A}}^{*}$ the Fenchel conjugate of ${\sf n}_{\mathscr{A}}\coloneqq\|\cdot\|_{L^{1}(T^{*}{\rm X};\mathscr{A})}\colon L^{1}(T^{*}{\rm X};\mathscr{A})\to[0,+\infty)$, i.e.
${\sf n}_{\mathscr{A}}^{*}(V)\coloneqq\sup_{\omega\in L^{1}(T^{*}{\rm X};\mathscr{A})}\big{(}V(\omega)-\|\omega\|_{L^{1}(T^{*}{\rm X};\mathscr{A})}\big{)}\quad\text{ for every }V\in L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}.$
Straightforward computations give that
${\sf n}_{\mathscr{A}}^{*}(V)=\left\{\begin{array}[]{ll}0\\ +\infty\end{array}\quad\begin{array}[]{ll}\text{ for every }V\in L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}\text{ with }\|V\|_{L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}}\leq 1,\\ \text{ for every }V\in L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}\text{ with }\|V\|_{L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}}>1.\end{array}\right.$ (4.7)
Theorem 4.8.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. Then it holds that ${\rm BV}_{\rm W}({\rm X};\mathscr{A})\subseteq{\rm BV}_{\rm H}({\rm X};\mathscr{A})$ and
$\|{\bf D}f\|_{*,\mathscr{A}}\leq|{\bf D}f|_{\mathscr{A}}({\rm X})\quad\text{ for every }f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A}).$
Proof.
Fix any $f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A})$. Since $\mathcal{E}_{\mathscr{A}}$ is convex and $L^{1}(\mathfrak{m})$-lower semicontinuous, it coincides with the double Fenchel conjugate ~$\tilde{\mathcal{E}}_{\mathscr{A}}^{**}\coloneqq(\tilde{\mathcal{E}}_{\mathscr{A}}^{*})^{*}$ of ~$\tilde{\mathcal{E}}_{\mathscr{A}}$. Therefore, using (4.6), [10, Theorem 5.1], (4.7) and Lemma 4.7, we obtain that
$\begin{split}\mathcal{E}_{\mathscr{A}}(f)&=\tilde{\mathcal{E}}_{\mathscr{A}}^{**}(f)=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg{(}\int fg\,{\mathrm{d}}\mathfrak{m}-\tilde{\mathcal{E}}_{\mathscr{A}}^{*}(g)\bigg{)}=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg{(}\int fg\,{\mathrm{d}}\mathfrak{m}-({\sf n}_{\mathscr{A}}\circ{\mathrm{d}}_{\mathscr{A}})^{*}(g)\bigg{)}\\ &=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg{(}\int fg\,{\mathrm{d}}\mathfrak{m}-\inf_{\begin{subarray}{c}V\in D({\mathrm{d}}_{\mathscr{A}}^{*}):\\ {\mathrm{d}}_{\mathscr{A}}^{*}V=g\end{subarray}}{\sf n}_{\mathscr{A}}^{*}(V)\bigg{)}=\sup_{V\in D({\mathrm{d}}_{\mathscr{A}}^{*})}\bigg{(}\int f\,{\mathrm{d}}_{\mathscr{A}}^{*}V\,{\mathrm{d}}\mathfrak{m}-{\sf n}_{\mathscr{A}}^{*}(V)\bigg{)}\\ &\leq\sup_{\begin{subarray}{c}b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}):\\ |b|_{\mathscr{A}}\leq 1\end{subarray}}\bigg{(}-\int f\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m}\bigg{)}=\sup_{\begin{subarray}{c}b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}):\\ |b|_{\mathscr{A}}\leq 1\end{subarray}}{\bf L}_{f}(b)({\rm X})=|{\bf D}f|_{\mathscr{A}}({\rm X}).\end{split}$
This gives that $f\in{\rm BV}_{\rm H}({\rm X};\mathscr{A})$ and $\|{\bf D}f\|_{*,\mathscr{A}}\leq|{\bf D}f|_{\mathscr{A}}({\rm X})$, thus proving the statement. ∎
Theorem 4.9.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Assume that $({\rm X},{\sf d})$ is either complete or a Radon space. Then it holds that
${\rm BV}_{\rm H}({\rm X})={\rm BV}_{\rm W}({\rm X})$
and we have that $|{\bf D}f|=|{\bf D}f|_{*}$ for every $f\in{\rm BV}_{\rm H}({\rm X})$.
Proof.
It follows from Lemma 3.15 and Theorem 4.8. ∎
Theorem 4.10.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Assume that $({\rm X},{\sf d})$ is complete. Let $\mathscr{A}$ be a good subalgebra of ${\rm LIP}_{b}({\rm X})$. Then it holds that
${\rm BV}_{\rm H}({\rm X};\mathscr{A})={\rm BV}_{\rm H}({\rm X})={\rm BV}_{\rm W}({\rm X};\mathscr{A})={\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X})).$
Moreover, we have that $\|{\bf D}f\|_{*,\mathscr{A}}=\|{\bf D}f\|_{*}=|{\bf D}f|_{\mathscr{A}}({\rm X})=|{\bf D}f|_{{\rm LIP}_{b}({\rm X})}({\rm X})$ and $|{\bf D}f|_{*}=|{\bf D}f|_{\mathscr{A}}=|{\bf D}f|_{{\rm LIP}_{b}({\rm X})}$ for every $f\in{\rm BV}_{\rm H}({\rm X})$.
Proof.
First of all, observe that for any unital separating subalgebra $\mathscr{A}\subseteq{\rm LIP}_{b}({\rm X})$ it holds that
$\begin{split}&{\rm BV}_{\rm W}({\rm X};\mathscr{A})\subseteq{\rm BV}_{\rm H}({\rm X};\mathscr{A})\subseteq{\rm BV}_{\rm H}({\rm X})\subseteq{\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X})),\\ &|{\bf D}f|_{{\rm LIP}_{b}({\rm X})}({\rm X})\leq\|{\bf D}f\|_{*}\leq\|{\bf D}f\|_{*,\mathscr{A}}\leq|{\bf D}f|_{\mathscr{A}}({\rm X})\quad\text{ for every }f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A}).\end{split}$ (4.8)
The above inclusions (from left to right) and the above inequalities (from right to left) follow from Theorem 4.8, (3.2) and Remark 3.16. If $\mathscr{A}$ is a good subalgebra, we know from Corollary 4.6 that ${\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X}))\subseteq{\rm BV}_{\rm W}({\rm X};\mathscr{A})$ and $|{\bf D}f|_{\mathscr{A}}\leq|{\bf D}f|_{{\rm LIP}_{b}({\rm X})}$ for every $f\in{\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X}))$. Therefore, the above inclusions and inequalities are all equalities. The statement follows. ∎
We conclude by formulating a partial converse of Theorem 4.10:
Proposition 4.11.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\mathscr{A}$ be a unital separating subalgebra of ${\rm LIP}_{\star}({\rm X})$. If ${\rm BV}_{\rm H}({\rm X};\mathscr{A})={\rm BV}_{\rm H}({\rm X})$ and $\|{\bf D}f\|_{*,\mathscr{A}}=\|{\bf D}f\|_{*}$ for every $f\in{\rm BV}_{\rm H}({\rm X};\mathscr{A})$, then $\mathscr{A}$ is a weakly good subalgebra: for every $x\in{\rm X}$ and for every $r>0$, there exist $(f_{n})_{n}\subseteq\mathscr{A}$ and $\rho\in L^{\infty}(\mathfrak{m})$ with $0\leq\rho\leq 1$ such that $\|f_{n}-{\sf d}_{x}\wedge r\|_{L^{1}(\mathfrak{m})}\to 0$ and ${\rm lip}_{a}(f_{n})\mathfrak{m}\rightharpoonup\rho\mathfrak{m}$ weakly.
Proof.
Fix any $x\in{\rm X}$ and $r>0$. Given that ${\sf d}_{x}\wedge r\in{\rm LIP}_{\star}({\rm X})\subseteq{\rm BV}_{\rm H}({\rm X})={\rm BV}_{\rm H}({\rm X};\mathscr{A})$ and $\|{\bf D}({\sf d}_{x}\wedge r)\|_{*,\mathscr{A}}=|{\bf D}({\sf d}_{x}\wedge r)|_{*}({\rm X})$, we can find $(f_{n})_{n}\subseteq\mathscr{A}$ such that $f_{n}\to{\sf d}_{x}\wedge r$ in $L^{1}(\mathfrak{m})$ and $\int{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}\to|{\bf D}({\sf d}_{x}\wedge r)|_{*}({\rm X})$. It follows from Remark 3.3 that ${\rm lip}_{a}(f_{n})\mathfrak{m}\rightharpoonup|{\bf D}({\sf d}_{x}\wedge r)|_{*}$ weakly. Since $|{\bf D}({\sf d}_{x}\wedge r)|_{*}\leq\mathfrak{m}$ as a direct consequence of the definition, we proved the statement. ∎
Appendix A The total variation measure
We report here a detailed version of the arguments in the proofs of [4, Lemmata 5.2 and 5.4] showing that they do not make use of the completeness nor separability of the ambient metric space, and also that they still work with our notion of asymptotic slope and for the approximating class being ${\rm LIP}_{\star}({\rm X})$ and not ${\rm LIP}_{loc}({\rm X})$.
Lemma A.1.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space and let $M,N\subseteq{\rm X}$ be open sets such that $\eta\coloneqq{\sf d}(M\setminus N,N\setminus M)>0$. Set
$\phi(x):=\frac{3}{\eta}\min\left\{\big{(}{\sf d}(x,N\setminus M)-\eta/3\big{)}^{+},\frac{\eta}{3}\right\}\quad\text{ for every }x\in M\cup N,$
and $H\coloneqq\phi^{-1}((0,1))$. Then $H$ is an open subset of $M\cap N$ with ${\sf d}(H,(M\setminus N)\cup(N\setminus M))>0$. Whenever $u\in{\rm LIP}_{\star}(M)$ and $v\in{\rm LIP}_{\star}(N)$, setting
$w(x)\coloneqq(1-\phi(x))u(x)+\phi(x)v(x)\quad\text{ for every }x\in M\cup N,$
(where we extended trivially $u$ and $v$ to the whole set $M\cup N$), we have that $w\in{\rm LIP}_{\star}(M\cup N)$ and moreover it holds:
$\int_{M\cup N}{\rm lip}_{a}(w)\,{\mathrm{d}}\mathfrak{m}\leq\int_{M}{\rm lip}_{a}(u)\,{\mathrm{d}}\mathfrak{m}+\int_{N}{\rm lip}_{a}(v)\,{\mathrm{d}}\mathfrak{m}+\frac{3}{\eta}\int_{\overline{H}}|u-v|\,{\mathrm{d}}\mathfrak{m}.$
$w\equiv u\text{ on a neighbourhood of $M\setminus N$},\quad w\equiv v\text{ on a neighbourhood of $N\setminus M$}.$
$\int_{M\cup N}|w-\sigma|\,{\mathrm{d}}\mathfrak{m}\leq\int_{M}|u-\sigma|\,{\mathrm{d}}\mathfrak{m}+\int_{N}|v-\sigma|\,{\mathrm{d}}\mathfrak{m}\quad\text{ for every }\sigma\in L^{1}(M\cup N).$
Proof.
It is clear that $\phi\in{\rm LIP}_{b}(M\cup N)$, that $\phi\equiv 1$ on $\{x\in M\cup N:{\sf d}(x,N\setminus M)>2\eta/3\}\supseteq M\setminus N$, that $\phi\equiv 0$ on $\{x\in M\cup N:{\sf d}(x,N\setminus M)<\eta/3\}\supseteq N\setminus M$, that $\phi(M\cup N)=[0,1]$, and that $\phi$ is $3/\eta$-Lipschitz continuous. The set $H\coloneqq\phi^{-1}((0,1))$ is therefore an open subset of $M\cap N$ and it satisfies ${\sf d}(H,(M\setminus N)\cup(N\setminus M))>0$: indeed $x\in H$ if and only if $\eta/3<{\sf d}(x,N\setminus M)<2\eta/3$, so that ${\sf d}(x,N\setminus M)>\eta/3$ and ${\sf d}(x,M/\setminus N)>\eta-2\eta/3=\eta/3$. Given $u$ and $v$ as in the statement, let $w=\phi u+(1-\phi)v$ in $M\cup N$ be as above. The convexity inequality for the asymptotic slope gives that ${\rm lip}_{a}(w)\leq\phi\,{\rm lip}_{a}(u)+(1-\phi){\rm lip}_{a}(v)+{\rm lip}_{a}(\phi)|u-v|$, so that $w\in{\rm LIP}_{\star}(M\cup N)$. By definition, it is clear that $w\equiv u$ in a neighbourhood of $M\setminus N$ and that $w\equiv v$ in a neighbourhood of $N\setminus M$. Writing $M\cup N=\operatorname{int}(\{\phi=0\})\cup\operatorname{int}(\{\phi=1\})\cup\overline{H}$, we have
$\int_{M\cup N}{\rm lip}_{a}(w)\,{\mathrm{d}}\mathfrak{m}$ $\leq\int_{M\cup N}\phi\,{\rm lip}_{a}(u)\,{\mathrm{d}}\mathfrak{m}+\int_{M\cup N}(1-\phi){\rm lip}_{a}(v)+\int_{M\cup N}{\rm lip}_{a}(\phi)|u-v|\,{\mathrm{d}}\mathfrak{m}$
$\leq\int_{M}{\rm lip}_{a}(u)\,{\mathrm{d}}\mathfrak{m}+\int_{N}{\rm lip}_{a}(v)\,{\mathrm{d}}\mathfrak{m}+\frac{3}{\eta}\int_{\overline{H}}|u-v|\,{\mathrm{d}}\mathfrak{m},$
where we have also used that ${\rm lip}_{a}(\phi)\equiv 0$ in the interior of sets of the form $\{\phi=c\}$, $c\in\{0,1\}$, and that $\phi\leq\mathbbm{1}_{M}$ and $1-\phi\leq\mathbbm{1}_{N}$. The last inequality in the statement follows from the latter fact and the triangle inequality, writing $\sigma=\phi\sigma+(1-\phi)\sigma$. ∎
Proposition A.2.
Let $({\rm X},{\sf d},\mathfrak{m})$ be a metric measure space. Let $\tau_{\sf d}$ be the topology induced by ${\sf d}$. For $f\in L^{1}(\mathfrak{m})$ we define the function $|{\bf D}f|_{*}:\tau_{\sf d}\to[0,+\infty]$ as
$|{\bf D}f|_{*}(U)\coloneqq\inf\bigg{\{}\liminf_{n\to\infty}\int_{U}{\rm lip}_{a}(f_{n})\,{\mathrm{d}}\mathfrak{m}\;\bigg{|}\;{\rm LIP}_{\star}(U)\ni f_{n}\to f|_{U}\text{ in }L^{1}(\mathfrak{m}|_{U})\bigg{\}}\quad\text{ for every }U\in\tau_{\sf d},$
with the convention that $|{\bf D}f|_{*}(\varnothing)\coloneqq 0$. Then $|{\bf D}f|_{*}$ satisfies the following properties:
  1. (1)
    If $A_{1},A_{2}\in\tau_{\sf d}$ and $A_{1}\subseteq A_{2}$, then $|{\bf D}f|_{*}(A_{1})\leq|{\bf D}f|_{*}(A_{2})$.
  2. (2)
    If $A_{1},A_{2}\in\tau_{\sf d}$, then $|{\bf D}f|_{*}(A_{1}\cup A_{2})\leq|{\bf D}f|_{*}(A_{1})+|{\bf D}f|_{*}(A_{2})$, with equality if $A_{1}$ and $A_{2}$ are disjoint.
  3. (3)
    If $(A_{n})_{n}\subseteq\tau_{\sf d}$ and $A_{n}\subseteq A_{n+1}$ for every $n\in\mathbb{N}$, then
    $\lim_{n\to\infty}|{\bf D}f|_{*}(A_{n})=|{\bf D}f|_{*}\left(\bigcup_{n=1}^{\infty}A_{n}\right).$ (A.1)
In particular the formula
$|{\bf D}f|_{*}(E)\coloneqq\inf\big{\{}|{\bf D}f|_{*}(U)\;\big{|}\;U\subseteq{\rm X}\text{ open, }E\subseteq U\big{\}}\quad\text{ for every }E\subseteq{\rm X},$ (A.2)
provides a $\sigma$-subadditive extension of $|{\bf D}f|_{*}$ whose collection of additive sets, in the sense of Carathéodory, contains $\mathcal{B}(X)$. Therefore $|{\bf D}f|_{*}\colon\mathcal{B}(X)\to[0,+\infty]$ is a Borel measure.
Proof.
The monotonicity in (1) and the equality in (2) in case $A_{1}$ and $A_{2}$ are disjoint are obvious. We proceed to prove (3), setting $A\coloneqq\bigcup_{n=1}^{\infty}A_{n}$.
Preparation to the proof of (3): by the monotonicity in (1), it is enough to show
$\sup_{n}|{\bf D}f|_{*}(A_{n})\geq|{\bf D}f|_{*}(A)$ (A.3)
so that we can assume w.l.o.g. that $\sup_{n}|{\bf D}f|_{*}(A_{n})<+\infty$. We can restrict ourselves to the case
${\sf d}(\overline{A_{n}},{\rm X}\setminus A_{n+1})>0\quad\text{ for every }n\in\mathbb{N}.$ (A.4)
To see this, it is enough to set
$A_{n}^{\prime}\coloneqq\left\{x\in{\rm X}:{\sf d}(X,{\rm X}\setminus A_{n})>\frac{1}{n}\right\}\in\tau_{\sf d},\quad\text{ for every }n\in\mathbb{N},$
so that $\bigcup_{n=1}^{\infty}A^{\prime}_{n}=A$, condition (A.4) is satisfied for $(A^{\prime}_{n})_{n}$ and, since $A_{n}^{\prime}\subseteq A_{n}$ for every $n\in\mathbb{N}$, by the monotonicity in (1), we also have $\sup_{n}|{\bf D}f|_{*}(A_{n})\geq\sup_{n}|{\bf D}f|_{*}(A^{\prime}_{n})$.
We therefore assume (A.4) to hold for $(A_{n})_{n}$ and we proceed to set
$C_{1}\coloneqq A_{1},\quad C_{2}\coloneqq A_{2},\quad C_{k}\coloneqq A_{k}\setminus\overline{A_{k-2}}\;\text{ for }k\geq 3.$
It is clear that $(C_{k})_{k}\subset\tau_{\sf d}$, ${\sf d}(C_{3k+i},C_{3h+i})>0$ for every $k,h\geq 0$, $h\neq k$, and every $i\in\{1,2,3\}$. By construction we also have that $\sum_{j=0}^{n}C_{3j+i}\subseteq A_{3n+i}$ for every $n\in\mathbb{N}$ and $i\in\{1,2,3\}$, so that,
$\sum_{k=1}^{\infty}|{\bf D}f|_{*}(C_{k})=\sum_{i=1}^{3}\sum_{j=0}^{\infty}|{\bf D}f|_{*}(C_{3j+i})\leq 3\sup_{n}|{\bf D}f|_{*}(A_{n})<+\infty,$
where we have also used the equality in (2) for disjoint sets. We fix $\varepsilon>0$ and we find an integer $\bar{k}\in\mathbb{N}$ such that $\sum_{k=\bar{k}}^{\infty}|{\bf D}f|_{*}(C_{k})\leq\varepsilon$. Point (3) will follow if we are able to construct a sequence $(f_{m})_{m}\subseteq{\rm LIP}_{\star}(A)$ such that
$|{\bf D}f|_{*}(A_{\bar{k}})+5\varepsilon\geq\liminf_{m\to\infty}\int_{A}{\rm lip}_{a}(f_{m})\,{\mathrm{d}}\mathfrak{m},\quad f_{m}\to f|_{A}\text{ in }L^{1}(\mathfrak{m}|_{A}),$ (A.5)
so that (A.3), and thus (3), follows. Notice that the sequence $(f_{m})_{m}$ (and many of the objects defined below) depends also on $\varepsilon$ but we do not stress this dependence. Let us set
$B_{-2}\coloneqq\varnothing,\quad B_{-1}\coloneqq\varnothing,\quad B_{0}\coloneqq A_{\bar{k}},\quad D_{0}\coloneqq A_{\bar{k}},\quad B_{h}\coloneqq A_{h+\bar{k}},\quad D_{h}\coloneqq C_{h+\bar{k}}\quad\text{ for all }h\in\mathbb{N}.$
Let $c_{h}>0$ and $H_{h}$ be the constant and the open set as in Lemma A.1 for the sets $B_{h}$ and $D_{h+1}$, for every $h\geq 0$; we also set $c_{-1}\coloneqq 0$ and $H_{-1}\coloneqq\varnothing$. Note that ${\sf d}(D_{h+1}\setminus B_{h},B_{h}\setminus D_{h+1})>0$ for every $h\geq 0$ by construction. To construct $f_{m}$ as above, we construct, for every $m\in\mathbb{N}$, sequences of functions $(\psi_{h}^{m})_{h\geq 0}$, $(f_{h}^{m})_{h\geq 0}$. Starting from those, we define a sequence $(\tilde{f}_{m})_{m}\subset{\rm LIP}_{loc}(A)$ satisfying (A.5) from which we will define the sought sequence $(f_{m})_{m}\subset{\rm LIP}_{\star}(A)$ satisfying (A.5) by a simple truncation argument.
Construction of $(\psi_{h}^{m})_{h}$: let $m\in\mathbb{N}$ and $h\geq 0$ be fixed and let us set
$\gamma_{h}^{m}\coloneqq\min\left\{\frac{1}{m2^{h}},\frac{\varepsilon}{2^{h+1}c_{h}},\frac{\varepsilon}{2^{h}c_{h-1}}\right\}>0.$
By definition of $|{\bf D}f|_{*}(D_{h})$, we can find a function $\psi_{h}^{m}\in{\rm LIP}_{\star}(D_{h})$ such that
$\int_{D_{h}}|\psi_{h}^{m}-f|\,{\mathrm{d}}\mathfrak{m}<\gamma_{h}^{m},\quad\int_{D_{h}}{\rm lip}_{a}(\psi_{h}^{m})\,{\mathrm{d}}\mathfrak{m}<|{\bf D}f|_{*}(D_{h})+\gamma_{h}^{m}.$
It follows that
$\int_{D_{h}}{\rm lip}_{a}(\psi_{h}^{m})\,{\mathrm{d}}\mathfrak{m}$ $\leq|{\bf D}f|_{*}(D_{h})+\frac{1}{m2^{h}},$ (A.6)
$\int_{D_{h}}|\psi_{h}^{m}-f|\,{\mathrm{d}}\mathfrak{m}$ $\leq\frac{1}{m2^{h}},$ (A.7)
$c_{h}\int_{\overline{H_{h}}}|\psi_{h}^{m}-f|\,{\mathrm{d}}\mathfrak{m}$ $\leq\frac{\varepsilon}{2^{h+1}},$ (A.8)
$c_{h-1}\int_{\overline{H_{h-1}}}|\psi_{h}^{m}-f|\,{\mathrm{d}}\mathfrak{m}$ $\leq\frac{\varepsilon}{2^{h}},$ (A.9)
where the latter two inequalities follow from the fact that $\overline{H_{h-1}}\subset B_{h-1}\cap D_{h}\subset D_{h}$ and $\overline{H_{h}}\subset B_{h}\cap D_{h+1}\subset D_{h}$.
Construction of $(f_{h}^{m})_{h}$: let $m\in\mathbb{N}$ be fixed; we construct by induction on $h$ a sequence of functions $(f_{h}^{m})_{h\geq-1}$ such that $f^{m}_{-1}:=0$ , $f^{m}_{0}:=\psi^{m}_{0}$ and satisfying, for every $h\geq 0$, the following conditions:
$f^{m}_{h}\equiv f^{m}_{h-1}\quad\text{ on }B_{h-2},\quad$ $f^{m}_{h}\equiv\psi^{m}_{h}\quad\text{ on }B_{h}\setminus\overline{B_{h-1}},\quad f^{m}_{h}\in{\rm LIP}_{\star}(B_{h}),$ (A.10)
$\int_{B_{h}}|f^{m}_{h}-f|\,{\mathrm{d}}\mathfrak{m}\leq\frac{1}{m}\left(2-\frac{1}{2^{h}}\right),$ (A.11)
$\int_{B_{h}}{\rm lip}_{a}(f^{m}_{h})\,{\mathrm{d}}\mathfrak{m}\leq$ $\int_{B_{h-1}}{\rm lip}_{a}(f^{m}_{h-1})\,{\mathrm{d}}\mathfrak{m}+\int_{D_{h}}{\rm lip}_{a}(\psi^{m}_{h})\,{\mathrm{d}}\mathfrak{m}+\frac{\varepsilon}{2^{h-1}}.$ (A.12)
Notice that the above conditions are satisfied for $h=0$: the first condition in (A.10) is void, the second and the third one are true by construction; (A.11) follows by (A.7); (A.12) is trivially true. We now show that the above three conditions being true for $h\geq 0$ implies they are true for $h+1$. We apply Lemma A.1 to the open sets $M\coloneqq B_{h}$ and $N\coloneqq D_{h+1}$ and to the functions $u\coloneqq f_{h}^{m}$ and $v\coloneqq\psi^{m}_{h+1}$. We get the existence of a function $f^{m}_{h+1}\in{\rm LIP}_{\star}(B_{h+1})$ satisfying
$\int_{B_{h+1}}{\rm lip}_{a}(f^{m}_{h+1})\,{\mathrm{d}}\mathfrak{m}\leq\int_{B_{h}}{\rm lip}_{a}(f_{h}^{m})\,{\mathrm{d}}\mathfrak{m}+\int_{D_{h+1}}{\rm lip}_{a}(\psi^{m}_{h+1})\,{\mathrm{d}}\mathfrak{m}+c_{h}\int_{\overline{H_{h}}}|f_{h}^{m}-\psi^{m}_{h+1}|\,{\mathrm{d}}\mathfrak{m},$ (A.13)
$f^{m}_{h+1}\equiv f_{h}^{m}\text{ on $B_{h}\setminus D_{h+1}\supseteq B_{h-1}$},\quad f^{m}_{h+1}\equiv\psi^{m}_{h+1}\text{ on $D_{h+1}\setminus B_{h}\supseteq B_{h+1}\setminus\overline{B_{h}}$},$ (A.14)
$\int_{B_{h+1}}|f^{m}_{h+1}-\sigma|\,{\mathrm{d}}\mathfrak{m}\leq\int_{B_{h}}|f_{h}^{m}-\sigma|\,{\mathrm{d}}\mathfrak{m}+\int_{D_{h+1}}|\psi^{m}_{h+1}-\sigma|\,{\mathrm{d}}\mathfrak{m}\quad\text{ for every }\sigma\in L^{1}(B_{h+1}).$ (A.15)
We see that (A.10) for $h+1$ follows from (A.14); choosing $\sigma:=f$ in (A.15) and using (A.11) and (A.7) we get (A.11) for $h+1$; finally (A.12) for $h+1$ follows from (A.13), (A.8) and (A.9) and the fact that by induction assumption $f_{h}^{m}\equiv\psi_{h}^{m}$ on $B_{h}\setminus\overline{B_{h-1}}\supseteq\overline{H_{h}}$.
Construction of $(f_{m})_{m}$ and proof of (3): for every $m\in\mathbb{N}$ we set
$\tilde{f}_{m}(x)\coloneqq f^{m}_{h}(x)\quad\text{ if }x\in B_{h-1},\quad h\geq 1.$
By the first equality in (A.10) each function $\tilde{f}_{m}$ is well defined and belongs to ${\rm LIP}_{loc}(A)$ for every $m\in\mathbb{N}$. Inequality (A.11) and the monotone convergence theorem give that $\int_{A}|\tilde{f}_{m}-f|\,{\mathrm{d}}\mathfrak{m}\leq 2/m$ for every $m\in\mathbb{N}$ so that $\tilde{f}_{m}\to f|_{A}\text{ in }L^{1}(\mathfrak{m}|_{A})$. Iterating (A.12) and using (A.6) we have that, for every $h\geq 1$, it holds
$\int_{B_{h}}{\rm lip}_{a}(f^{m}_{h})\,{\mathrm{d}}\mathfrak{m}$ $\leq\int_{B_{h-1}}{\rm lip}_{a}(f^{m}_{h-1})\,{\mathrm{d}}\mathfrak{m}+\int_{D_{h}}{\rm lip}_{a}(\psi^{m}_{h})\,{\mathrm{d}}\mathfrak{m}+\frac{\varepsilon}{2^{h-1}}$
$\leq\int_{B_{h-1}}{\rm lip}_{a}(f^{m}_{h-1})\,{\mathrm{d}}\mathfrak{m}+|{\bf D}f|_{*}(D_{h})+\frac{1}{2^{h}}\left(\frac{1}{m}+2\varepsilon\right)$
$\leq\sum_{i=0}^{h}|{\bf D}f|_{*}(D_{h})+\left(\frac{1}{m}+2\varepsilon\right)\sum_{i=0}^{h}\frac{1}{2^{h}}$
$\leq|{\bf D}f|_{*}(A_{\bar{k}})+\sum_{k=\bar{k}}^{h+\bar{k}}|{\bf D}f|_{*}(C_{k})+\left(\frac{1}{m}+2\varepsilon\right)\sum_{i=0}^{h}\frac{1}{2^{h}}$
$\leq|{\bf D}f|_{*}(A_{\bar{k}})+\varepsilon+\left(\frac{1}{m}+2\varepsilon\right)\sum_{i=0}^{h}\frac{1}{2^{h}}.$
Therefore, the monotone convergence theorem gives that
$\int_{A}{\rm lip}_{a}(\tilde{f}_{m})\,{\mathrm{d}}\mathfrak{m}=\lim_{h\to\infty}\int_{B_{h}}{\rm lip}_{a}(f_{h}^{m})\,{\mathrm{d}}\mathfrak{m}\leq|{\bf D}f|_{*}(A_{\bar{k}})+5\varepsilon+\frac{2}{m},$
which shows that $(\tilde{f}_{m})_{m}\subset{\rm LIP}_{loc}(A)$ satisfies (A.5). Finally we set
$f_{m}\coloneqq\tilde{f}_{m}\wedge m\in{\rm LIP}_{\star}(A)\quad\text{ for every }m\in\mathbb{N}.$
We see that $f_{m}\to f|_{A}\text{ in }L^{1}(\mathfrak{m}|_{A})$ and, since ${\rm lip}_{a}(f_{m})\leq{\rm lip}_{a}(\tilde{f}_{m})$ for every $m\in\mathbb{N}$, we also see that $(f_{m})_{m}\subset{\rm LIP}_{\star}(A)$ satisfies the first part of (A.5). This concludes the proof of (3).
Proof of (2): it is enough to show
$|{\bf D}f|_{*}(A_{1}^{\prime}\cup A_{2}^{\prime})\leq|{\bf D}f|_{*}(A_{1})+|{\bf D}f|_{*}(A_{2})\quad\text{ if }A_{i}^{\prime}\in\tau_{\sf d}\text{ is s.t.~{}}{\sf d}(A_{i}^{\prime},{\rm X}\setminus A_{i})>0,\,i=1,2.$ (A.16)
Indeed, it is enough to define
$A_{i}^{n}\coloneqq\left\{x\in{\rm X}:{\sf d}(x,{\rm X}\setminus A_{i})>1/n\right\}\in\tau_{\sf d},\quad\text{ for everr }i=1,2,\,n\in\mathbb{N},$
so that, if (A.16) is true for $A_{i}^{\prime}\coloneqq A_{i}^{n}$, $i=1,2$, $n\in\mathbb{N}$, then (3) gives that $\sup_{n}|{\bf D}f|_{*}(A_{1}^{n}\cup A_{2}^{n})=|{\bf D}f|_{*}(A_{1}\cup A_{2})$ and we obtain (2). To show (A.16) we take sequences $(g^{i}_{h})_{h}\subset{\rm LIP}_{\star}(A_{i})$ satisfying
$\lim_{h\to\infty}\int_{A_{i}}{\rm lip}_{a}(g^{i}_{h})\,{\mathrm{d}}\mathfrak{m}=|{\bf D}f|_{*}(A_{i}),\quad g^{i}_{h}\to f\text{ in }L^{1}(\mathfrak{m}|_{A_{i}}),\quad i=1,2,$ (A.17)
and we apply Lemma A.1 with $M\coloneqq(A_{1}^{\prime}\cup A_{2}^{\prime})\cap A_{1}$, $N\coloneqq(A_{1}^{\prime}\cup A_{2}^{\prime})\cap A_{2}$, and, for every $h\in N$, to the functions $u\coloneqq g_{h}^{1}$, $v\coloneqq g_{h}^{2}$. We obtain the existence of a positive constant $c>0$ and, for every $h\in\mathbb{N}$, of a function $g_{h}\in{\rm LIP}_{\star}(A_{1}^{\prime}\cup A_{2}^{\prime})$ satisfying
$\int_{A_{1}^{\prime}\cup A_{2}^{\prime}}{\rm lip}_{a}(g_{h})\,{\mathrm{d}}\mathfrak{m}\leq\int_{A_{1}}{\rm lip}_{a}(g_{h}^{1})\,{\mathrm{d}}\mathfrak{m}+\int_{A_{2}}{\rm lip}_{a}(g_{h}^{2})\,{\mathrm{d}}\mathfrak{m}+c\int_{A_{1}\cap A_{2}}|g_{h}^{1}-g_{h}^{2}|\,{\mathrm{d}}\mathfrak{m}.$
Passing to the limit as $h\to\infty$ and using (A.17), we get (A.16).
Finally, the given properties (1), (2), (3) imply that (A.2) defines an extension of $|{\bf D}f|_{*}$ to the family of all subsets of ${\rm X}$ which is an outer measure on ${\rm X}$ [9, Definition 1.11.1]; therefore its restriction to the $\sigma$-algebra of its Carathéodory-measurable subsets $\mathcal{A}$ is a $\sigma$-additive measure [9, Theorem 1.11.4(iii)]. Finally, we observe that $\mathcal{B}({\rm X})\subseteq\mathcal{A}$ as a consequence of [9, Theorem 1.11.10, Theorem 7.14.29] and (2). ∎